### v1

# Reflectance Prediction-based Knowledge Distillation for Robust 3D Object Detection in Compressed Point Clouds

Abstract
Regarding intelligent transportation systems for vehicle networking, low-bitrate transmission via lossy point cloud compression is vital for facilitating real-time collaborative perception among vehicles with restricted bandwidth. In existing compression transmission systems, the sender lossily compresses point coordinates and reflectance to generate a transmission code stream, which faces transmission burdens from reflectance encoding and limited detection robustness due to information loss. To address these issues, this paper proposes a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD). We compress point coordinates while discarding reflectance during low-bitrate transmission, and feed the decoded non-reflectance compressed point clouds into a student detector. The discarded reflectance is then reconstructed by a geometry-based reflectance prediction (RP) module within the student detector for precise detection. A teacher detector with the same structure as student detector is designed for performing reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) from raw to compressed point clouds. Our RPKD framework jointly trains detectors on both raw and compressed point clouds to improve the student detector’s robustness. Experimental results on the KITTI dataset and Waymo Open Dataset demonstrate that our method can boost detection accuracy for compressed point clouds across multiple code rates. Notably, at a low code rate of 2.146 Bpp on the KITTI dataset, our RPKD-PV achieves the highest mAP of 73.6, outperforming existing detection methods with the PV-RCNN baseline.
Index Terms:
Compressed Point Clouds, 3D Object Detection, Knowledge Distillation, Reflectance Prediction.I Introduction
LiDAR point clouds, composed of sparse points with both geometric (coordinates) and attribute (reflectance) information about object surfaces, are vital for predicting the categories, positions, and bounding boxes of 3D objects. In cooperative perception among networked vehicles, local vehicles (receivers) need to rapidly acquire LiDAR point clouds captured by neighboring vehicles (senders) to enable an expanded perception range and accurate object detection [1, 2, 3, 4]. Raw point clouds are extremely voluminous and require a high transmission bandwidth, making it difficult to realize real-time cooperative perception. To this end, point cloud compression reduces the amount of transmitted data, improving data acquisition efficiency at the receiver. However, compressed point clouds generated by existing methods often lose some geometric coordinates and reflectance information, leading to suboptimal performance in 3D object detection. Therefore, enhancing the 3D detector’s robustness for compressed point clouds is of great research significance.
Existing 3D object detection methods in compressed point clouds can be classified into two categories based on their reflectance processing approaches: 3D object detection with or without reflectance encoding. For 3D object detection with reflectance encoding, the geometry-based point cloud compression (G-PCC) [5] separately compresses the coordinates and reflectance of voxelized point clouds to form geometry and attribute bitstreams for reconstruction and detection tasks, as shown in Fig. 1(a). Building on this, Unicorn [6, 7] introduces a learning-based compression framework that utilizes multiscale sparse tensors from single-frame voxelized point clouds and lower-scale priors from multi-frame point clouds to improve compression and reconstruction performance. However, these methods face a significant burden from the attribute bitstream during networking transmission, limiting transmission efficiency. To solve this issue, in 3D object detection without reflectance encoding, existing compression methods based on voxelized point clouds and deep learning [8, 9, 10, 11, 12, 13] typically focus on compressing geometric coordinates while discarding reflectance to elevate transmission efficiency, as presented in Fig. 1(b). As an illustration, PCC-S [9] employs a geometric perception module to improve the learning of sibling and ancestor contexts in an octree structure for excellent geometric coordinates compression and reconstruction, during which the decoded data is termed non-reflectance compressed point clouds. However, lossy compression by these methods causes substantial information loss in low-bitrate transmission, severely impacting 3D object detection accuracy. To address these challenges, this paper aims to improve the transmission efficiency of lossy compressed point clouds and develop robust compressed point-cloud detectors, supported by a series of detailed studies.
Fig. 1(c) shows a baseline of our method. It takes accurate object predictions from raw point clouds as teacher knowledge for distillation during detection training. In this setup, the teacher model transfers its knowledge to compressed point-cloud detectors through response-based detection knowledge distillation (DKD). As shown in Fig. 2, SparseKD-PV outperforms PV-RCNN by an average of 1.11% in mAP across multiple data types, whereas it still exhibits a significant detection performance gap compared to PV-RCNN’s mAP of 75.56 on raw data. In general, LiDAR point cloud reflectance provides crucial cues for identifying 3D objects, thus improving the reliability of 3D object detection. In SparseKD-PV, the decision to discard reflectance information to enhance transmission efficiency is a key factor contributing to the suboptimal performance. While methods like Unicorn [7] reconstruct reflectance via a residual-based attribute bitstream using a DNN network, and PCC-S [9] does not transmit the attribute bitstream but needs alternative strategies to recover reflectance, we argue that relying solely on non-reflectance compressed point clouds for detection misses important reflectance features. To address this limitation, we propose a dedicated reflectance prediction (RP) module at the receiver, which utilizes the geometric information of compressed point clouds and reflectance priors from raw point clouds. Additionally, to further enhance the accuracy of reflectance prediction, we introduce response-based reflectance knowledge distillation (RKD), which effectively guides the RP module during learning process, improving detection performance while maintaining efficient data transmission.
Incorporating the previous analysis, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds, as depicted in Fig. 1(d). Since compressed and raw point clouds lack direct point-to-point correspondence, we design a reflectance cross-match (RCM) module to assign reflectance labels from raw point-cloud voxels to compressed points. To construct suitable knowledge for reflectance distillation, we introduce a comparable reflectance inter-match (RIM) module to generate reflectance labels for the raw point-cloud detector. Besides, the proposed reflectance prediction (RP) module comprehensively extracts geometric features of voxelized point clouds, achieving reflectance reconstruction at the receiver. Through our reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD), rich teacher knowledge is transferred from raw to compressed point clouds for more precise reflectance reconstruction and robust 3D object detection. As shown in Fig. 2, RPKD-PV significantly increases the mAP values across multiple data types, which demonstrates the effectiveness of our RPKD framework for non-reflectance compressed point clouds.
In summary, the contributions of this work are as follows.
We propose an RCM module to generate reflectance labels for compressed points based on the corresponding raw point-cloud voxels.
At the receiver, we design a geometry-based RP module to reconstruct discarded reflectance, thereby improving detection performance for non-reflectance compressed point clouds.
We introduce two response-based distillation constraints, namely RKD and DKD, to transfer supervisory knowledge from raw to compressed point-cloud detectors, further enhancing detection robustness for compressed point clouds.
The remainder of the paper is organized as follows. Section II reviews existing studies in terms of 3D object detection, point cloud compression, and knowledge distillation, highlighting relevant technologies. In Section III, we present the overall architecture of a 3D object detection framework with reflectance prediction-based knowledge distillation and elaborate on its key components. Section IV demonstrates the effectiveness of our method on the KITTI dataset and Waymo Open Dataset and includes ablation studies to evaluate the effect of each component. Finally, Section V concludes this paper.
II Related Work
II-A 3D Object Detection
In 3D object detection for raw point clouds, existing methods can be sorted into two classes according to the data format of network input: point-based [14, 15, 16, 17, 18, 19, 20, 21] and voxel-based [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] methods. Point-based detectors extract hierarchical point-wise features by various sampling and aggregation methods to predict 3D objects. In contrast, voxel-based detectors utilize 3D sparse convolution to extract voxel features, and these features are projected onto the Bird’s Eye View (BEV) for 3D object detection. In 3D object detection for compressed point clouds, those generated by range image-based [32, 33, 34, 35] and voxel-based [8, 9] compression frameworks are input into compressed point-cloud detectors for recognition and localization. These methods generally rely on raw point clouds for pre-training 3D detectors and assess detection performance on lossy compressed point clouds across various code rates using the pre-trained models, as shown in Fig. 1(a) and 1(b). However, they lack effective robustness measures against information loss and reconstruction errors caused by lossy point cloud compression, leading to a significant decline in detection accuracy.
For robust 3D object detection in compressed point clouds, this paper addresses three key aspects. First, voxelized representation of compressed point clouds helps mitigate the negative effects of lossy compression. Thus, we develop robust compressed point-cloud detectors on voxel-based detection frameworks. Second, pre-training 3D detectors solely on raw point clouds is insufficient for adapting to the characteristics of compressed point clouds. To improve detection performance, our 3D detectors are trained on both raw and compressed point clouds. Finally, with a two-source joint training scheme, we use voxel-based two-stage detectors like PV-RCNN [22] and Voxel-RCNN [23] as backbone networks and transfer detection knowledge of the first-stage proposals from raw to compressed point-cloud detectors, thereby enhancing detection robustness for compressed point clouds.
II-B Point Cloud Compression
LiDAR point cloud compression methods can be categorized into three classes based on data representation: point-based [36, 37, 38], image-based [32, 33, 34, 39, 40, 35, 41, 42, 43], and voxel-based [5, 6, 7, 8, 9, 10, 11, 12, 13, 44] methods. Point-based methods utilize flexible point sampling and feature extraction to generate point-wise features, achieving point cloud compression and reconstruction. However, these methods are confronted with uneven 3D object loss and noise-point interference due to the lack of spatial structure and context information during lossy compression. Image-based methods project point clouds onto 2D images for quantization and encoding, but they introduce more severe three-dimensional spatial distortion and information loss compared to voxel-based methods. By analyzing octree or multi-scale sparse tensors, voxel-based methods enhance compression performance via context feature learning, multi-scale joint optimization, and redundant encoding reduction. Among these compression methods, only G-PCC [5] and Unicorn [7] incorporate reflectance encoding modules, as reflectance encoding presents challenges in structural design, end-to-end optimization, and data-flow overhead. To solve this problem, we design an RP module at the receiver to supplement reflectance information for detection training on non-reflectance compressed point clouds generated by voxel-based methods. Our approach effectively alleviates the design complexity of LiDAR point cloud compression and provides an end-to-end optimization strategy for reflectance prediction at the receiver.
II-C Knowledge Distillation
Knowledge distillation methods for 3D object detection can be divided into three categories according to their research objectives: model compression distillation [45, 46, 47, 48, 49, 50], detection improvement distillation [51, 52], and weak-data robustness distillation [53, 54, 55]. In model compression distillation, SparseKD [45] introduces response-based constraints using Logit and Label KD to guide lightweight student models, achieving accurate and efficient 3D object detection. In detection improvement distillation, SPNet [51] assigns category values to the interior points of ground truth (GT) boxes, while X-Ray [52] addresses object point loss caused by sparsity and occlusion utilizing tracking information from consecutive-frame point clouds. Both methods augment data representations with strong semantic priors, providing valuable supervisory knowledge for distillation. In weak-data robustness distillation, student models employ low-quality point clouds from rainy conditions [55] or low-resolution sensor data [53, 54], learning complete and accurate knowledge from raw point clouds to improve detection robustness in weak data scenarios. Inspired by these weak-data robustness distillation methods, our RPKD framework treats non-reflectance compressed point clouds involved in vehicle networking data transmission as weak data for student models and raw point clouds as guiding data for teacher models, aiming to develop robust compressed point-cloud detectors.
III Proposed Approach
III-A Overview of our Framework
During point cloud compression, reflectance encoding faces significant challenges in both end-to-end optimization and data-flow overhead. According to the separable compression mechanism of G-PCC [5], existing methods typically disregard reflectance compression and focus solely on optimizing geometric coordinate encoding to achieve efficient transmission and accurate detection. As a result, decoded compressed point clouds often lack reflectance information and suffer from point number reduction and reconstruction errors, which limit 3D object detection performance. To address this issue, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD), as shown in Fig. 3, to obtain more robust and precise 3D detectors against non-reflectance compressed point clouds.
First, our reflectance cross-match (RCM) module correlates lossy compressed points with neighboring raw point-cloud voxels. The nearest voxel’s average reflectance is assigned as the reflectance label for the current compressed point. Similarly, the reflectance inter-match (RIM) module assigns a voxel’s average reflectance as the reflectance label for corresponding raw points. Next, the proposed reflectance prediction (RP) module generates point-wise geometric features using 3D sparse convolution and voxel set extraction to predict reflectance. The predicted point-wise reflectance is then incorporated into corresponding non-reflectance voxels for subsequent detection tasks. Finally, we treat non-reflectance compressed point clouds as weak data and introduce two response-based knowledge distillation constraints: reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD). Specifically, RKD transfers reflectance prediction knowledge from raw point clouds via the RP module, while DKD conveys first-stage proposal knowledge from the two-stage 3D detector for raw point clouds. In what follows, we detail the proposed framework.
III-B Reflectance Label Generation
Due to point number reduction and reconstruction errors from lossy compression, compressed points do not correspond one-to-one with raw points. For compressed point reflectance prediction, we introduce an RCM module based on spatial relationships to assign reflectance labels from raw point-cloud voxels to corresponding compressed points. These labels, used as prior knowledge for reflectance prediction, are generated frame by frame during detection training and do not need to be compressed or transmitted. Additionally, to transfer effective knowledge during reflectance distillation, we also design a similar RIM module to acquire reflectance labels for raw points. The principles of these modules are outlined below.
Reflectance Cross-match Module. For geometric coordinate encoding, voxel-based compression methods [6, 9] are often employed to obtain non-reflectance compressed point clouds . Before training, with each point’s reflectance assigned to 0, the compressed point clouds fed into our detection networks are . Raw point clouds are represented as . These point clouds are voxelized to the same scale, generating compressed point-cloud voxels and raw point-cloud voxels . When using , and as inputs, the principle of our RCM module can be described as follows.
As depicted in Fig. 4, there are two spatial matching strategies between compressed points and raw point-cloud voxels: center matching and surrounding matching. Specifically, the voxel corresponding to compressed point coordinates is denoted as , and the surrounding non-empty voxels within a 1-voxel matching range are . The initial reflectance labels for compressed points are defined as:
where represents the average reflectance of in center matching, and is the nearest voxel’s average reflectance in surrounding matching. However, when multiple compressed points exist within the same compressed point-cloud voxel , their reflectance labels may come from different raw point-cloud voxels, resulting in significant numerical discrepancies that negatively impact compressed point reflectance prediction. To address this issue, we further calculate the average reflectance labels for all points within each compressed point-cloud voxel . The final reflectance labels for compressed points are thus expressed as:
where denotes the reflectance label for each point within the current compressed point-cloud voxel , and represents the averaging operation applied to these reflectance labels.
Reflectance Inner-match Module. In the RPKD framework, it is crucial to establish suitable reflectance prediction knowledge from raw point clouds to guide the prediction process for compressed point-cloud reflectance. Specifically, we treat the reflectance predictions for compressed and raw point clouds as a student-teacher pair, where both sets of labels need to maintain spatial correspondence and information similarity. In the RCM module, the reflectance labels for compressed points are derived from the corresponding raw point-cloud voxels. Therefore, our RIM module avoids directly using raw point reflectance and instead assigns the average reflectance of raw point-cloud voxels to corresponding points as their reflectance labels. Since raw points exactly correspond to their voxels, our RIM module employs center matching. With raw point coordinates and voxels as inputs, the raw point reflectance labels are computed by the following formula:
where is the voxel corresponding to raw point coordinates , and denotes the average reflectance of this voxel.
III-C Reflectance Prediction Module
Concerning the reflectance reconstruction of compressed point clouds, Unicorn [7] utilizes current-scale reflectance residual features and low-scale reflectance priors through a DNN module to predict reflectance. In contrast, our RP module exploits the geometric features of non-reflectance compressed point clouds and the reflectance priors of raw point clouds to achieve precise reflectance reconstruction at the receiver. This innovative design not only reduces the transmission rate in compression transmission systems but also enhances the robustness of 3D detectors against non-reflectance compressed point clouds. As shown in Fig. 3, the structure and principle of our RP module are detailed as follows.
RP Module for Compressed Point Clouds. In the compressed point-cloud detector, the RP module takes non-reflectance compressed point-cloud voxels as input. To accurately predict reflectance for compressed points, the RP module employs 3D sparse convolution to extract voxel features and voxel set extraction to refine point-wise features. We define the geometric features of compressed points as:
where denotes a sparse tensor transformation module, is the 3D sparse convolution with a downsampling size (i.e., the initial size) in the voxel-wise basic backbone, and represents the voxel set extraction corresponding to . The voxel set extraction structure in our RP module is identical to that in PV-RCNN [22], but the number of sampling points is equal to the number of compressed points in the current frame. Since the reflectance is a single-channel floating-point scalar in the range of , we use two fully connected (FC) layers as a prediction head to generate compressed point reflectance, denoted as . The predicted reflectance of compressed points, , is then incorporated into the corresponding compressed point-cloud voxels based on their spatial relationship for subsequent training and detection.
We screen the compressed point reflectance labels with values greater than zero, and define their indices as . We apply Mean Squared Error Loss (MSELoss) to constrain compressed point reflectance prediction, and the corresponding loss function is expressed as follows:
where represents the number of compressed point reflectance labels with values greater than zero.
RP Module for Raw Point Clouds. In 3D object detection knowledge distillation, maintaining similar network structures between teacher and student detectors benefits the efficient transfer of teacher knowledge. Therefore, in this study, we design an analogous RP module in the raw point-cloud detector to predict raw point reflectance. Non-reflectance raw point-cloud voxels are fed into this module, and the geometric features of raw points are represented as:
Subsequently, we predict raw point reflectance and incorporate it into corresponding voxels to obtain raw point-cloud voxels with the recovered reflectance .
We screen the raw point reflectance labels that are greater than zero and define their indices as . The loss function for raw point reflectance prediction is defined as:
where represents the number of raw point reflectance labels with values greater than zero.
III-D Reflectance Prediction-based Knowledge Distillation
In lossy compression, non-reflectance compressed point clouds undergo point number reduction and reconstruction errors, leading to incomplete and biased geometric information. If only these compressed point clouds are used for reflectance prediction and object detection, it is difficult to match the detection accuracy of raw point clouds. Drawing from the weak-data robustness research [53, 54, 55] in 3D object detection knowledge distillation, we treat the non-reflectance compressed point clouds as weak students, enabling them to learn accurate and complete knowledge from raw point clouds to improve detection performance for compressed point clouds. Specifically, we propose two components: RKD, which transfers reflectance prediction knowledge from our RP module, and DKD, which refines the first-stage proposals in the two-stage 3D object detector, as depicted in Fig. 3. Both proposed distillation methods fall under the response-based knowledge distillation category, and their principles are outlined as follows.
Reflectance Knowledge Distillation. Building upon the RP module’s prediction outputs, we obtain the compressed point reflectance and raw point reflectance . Since they do not have a one-to-one correspondence, we map the recovered teacher reflectance from raw point-cloud voxels to corresponding compressed points. The mapped teacher reflectance of each compressed point serves as a soft label of RKD. According to the spatial relationship in the RCM module, the raw point-cloud voxel corresponding to compressed point coordinates is denoted as , and the RKD soft labels are obtained by the following formula:
where represents the mapped reflectance value from a raw point-cloud voxel to corresponding compressed point coordinates , and is the number of compressed points that can be successfully matched with raw point-cloud voxels in the RCM module. We also adopt MSELoss as the loss function to constrain the reflectance distillation process, and the RKD loss function is formulated as
where denotes screening foreground compression points during loss computation, and is the number of foreground compression points.
Detection Knowledge Distillation. In Fig. 3, the two-stage 3D detector transfers key features from the first-stage network to the second-stage detector for detection refinement. Additionally, the first-stage dense proposals passed to the second stage inherently carry valuable detection knowledge for distillation. DKD combines the compressed point-cloud proposals and raw point-cloud proposals to form a student-teacher pair for response-based knowledge distillation. In these proposals, and represent the sigmoid-activated classification scores and bounding boxes, respectively. For distillation constraints, we apply Logit KD proposed in SparseKD [45] to design the distillation loss. Therefore, the DKD loss function can be expressed as:
where refers to the MSELoss of classification scores , represents the Smooth-L1 loss of bounding boxes , and is the number of foreground objects in the first-stage proposals.
III-E Loss Function
In this study, we pre-train separate detectors for the compressed and raw point clouds and initialize the training parameters of our RPKD framework using these pre-trained models. Therefore, the compressed point-cloud detection loss and the raw point-cloud detection loss are defined as follows:
where , , and represent the reflectance prediction, classification, bounding box and orientation losses of the first-stage detector, respectively. The corresponding weight coefficients are , , and . denotes the second-stage detector loss with a weight coefficient of 1. On this basis, the overall loss function of our RPKD framework is formulated as follows:
where , and are the classification, bounding box and orientation losses of the first-stage detector based on Label KD in SparseKD [45]. The weight coefficients for the distillation losses are and .
IV Experiments
For 3D object detection in compressed point clouds, we utilized LiDAR point clouds from the KITTI dataset and Waymo Open Dataset (WOD) as raw point-cloud inputs for the compression network. Subsequently, we compressed and reconstructed the geometric coordinates of these point clouds at multiple code rates through deep learning-based PCC-S, resulting in non-reflectance compressed point clouds. The proposed RPKD framework was trained on both compressed and raw point clouds, using only compressed point clouds for inference. Extensive experiments demonstrate that our method significantly enhances detection robustness for non-reflectance compressed point clouds generated by voxel-based compression. The following sections provide a comprehensive explanation of the implementation details, result comparisons, and ablation analysis.
| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
| Voxel-RCNN | 92.55 | 84.65 | 82.41 | 66.07 | 59.30 | 53.73 | 92.81 | 73.55 | 68.97 | 74.89 | |
| RP-Voxel(T3) | 92.18 | 84.12 | 81.75 | 62.04 | 56.11 | 51.05 | 93.67 | 72.87 | 68.41 | 73.58 | |
| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
| RPKD-PV(Ours) | 92.02 | 83.13 | 80.21 | 65.72 | 58.68 | 53.72 | 90.39 | 68.63 | 63.92 | 72.94 | |
| Voxel-RCNN(Baseline) | 92.27 | 82.79 | 80.17 | 63.64 | 56.21 | 51.21 | 90.69 | 71.35 | 66.92 | 72.81 | |
| RPKD-Voxel(Ours) | 92.06 | 82.91 | 79.98 | 70.40 | 64.14 | 57.63 | 92.26 | 70.31 | 65.84 | 75.06 | |
| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
| RPKD-PV(Ours) | 91.38 | 82.95 | 79.96 | 70.93 | 62.81 | 56.39 | 89.21 | 66.68 | 62.09 | 73.60 | |
| Voxel-RCNN(Baseline) | 92.03 | 82.72 | 80.16 | 61.49 | 54.77 | 49.30 | 88.27 | 67.99 | 63.97 | 71.19 | |
| RPKD-Voxel(Ours) | 91.78 | 83.00 | 80.30 | 70.33 | 63.36 | 57.17 | 92.12 | 68.10 | 63.54 | 74.41 | |
| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
| RPKD-PV(Ours) | 91.53 | 82.52 | 79.45 | 63.24 | 57.24 | 51.90 | 85.51 | 61.70 | 57.68 | 70.09 | |
| Voxel-RCNN(Baseline) | 90.96 | 81.37 | 78.73 | 60.10 | 51.66 | 45.88 | 83.05 | 62.86 | 58.62 | 68.14 | |
| RPKD-Voxel(Ours) | 90.67 | 81.87 | 78.95 | 65.95 | 59.02 | 53.54 | 87.79 | 64.80 | 60.35 | 71.44 | |
IV-A Implementation Details
IV-A1 Datasets and Hardware
For the KITTI dataset, we used 7481 labeled training samples as raw point clouds, dividing them into a training split with 3712 samples and a validation split with 3769 samples. Firstly, the PCC-S network employed the training split of raw point clouds for compression learning and reconstructed 7481 frames of point clouds through the trained compression model at octree levels of 12, 11, and 10, respectively. Regarding the training configuration, we trained the PCC-S network for 20 epochs on a single RTX 8000 GPU with a batch size of 2 and a maximum octree level of 12. Then, we established independent models for each code rate and conducted joint training on both the compressed and raw point clouds to build robust 3D detectors for compressed point clouds. These detectors were trained for 80 epochs on 4 RTX 4090 GPUs with a batch size of 16.
WOD is a large-scale detection dataset composed of 798 training sequences and 202 validation sequences. Due to the limited training efficiency of PCC-S and the computing power of a single RTX 8000 GPU, we only selected 20 training sequences and 10 validation sequences for research. Thus, the raw point clouds consisted of a training split with 3966 samples and a validation split with 1967 samples. Using the same training configuration as KITTI, we obtained 5933 frames of compressed point clouds by PCC-S on WOD at octree levels 12, 11, and 10, respectively. Additionally, the detectors of our RPKD framework were trained for 30 epochs on 4 RTX 4090 GPUs with a batch size of 8.
IV-A2 Training Parameters
In the RCM module, we matched compressed points with raw point-cloud voxels within a 1-voxel matching range, corresponding to spatial voxels. During pre-training, the compressed and raw point-cloud detectors shared the same loss weights, and these loss weights were set to . In the knowledge distillation stage, we designed two distillation methods, RKD and DKD, whose loss weights were set to . Specifically, the distillation constraints for DKD consisted of classification distillation loss and bounding box distillation loss , and their loss weights were . For dataset configuration, we trained compressed point-cloud detectors using the standard parameters of public datasets and expanded single-source data to two-source data for joint processing.
IV-A3 Evaluation Metrics
We evaluated the robustness of our RPKD framework for 3D object detection on compressed point clouds using the KITTI and WOD validation splits. The primary metric for KITTI was average precision (AP) at 40 recall positions with 3D IoU thresholds of 0.7, 0.5, and 0.5 for cars, pedestrians, and cyclists, respectively. For WOD, the metrics were AP and average precision weighted by heading (APH). To compare performance across multiple code rates and methods, we used mean Average Precision (mAP) and mean APH (mAPH) across categories and difficulty levels. The Bit-per-point (Bpp) metric represents the average number of bits required to encode each point, indicating the code rate of compressed point clouds.
IV-A4 Controlled Experiment Design
Firstly, in terms of data types, raw data, PCC-S-C12, PCC-S-C11, and PCC-S-C10 represented raw point clouds, the compressed point clouds of octree levels 12, 11, and 10, respectively. Then, the detection performance of PV-RCNN at each code rate served as the baseline, and the designed control experiments involved three detection methods, namely SparseKD-PV, RPV-KD, and RPKD-PV. In SparseKD-PV, with the teacher detector of PV-RCNN (T1) on raw data, we utilized Logit KD and Label KD for distillation learning and performance evaluation of compressed point-cloud detectors. RPV-KD and SparseKD-PV share the same teacher model and distillation methods, while the key difference is that RPV-KD mapped raw point reflectance into corresponding compressed points based on the matching principle of our RCM module (Section III-B) to simulate compressed point clouds with decoded reflectance. RPKD-PV denotes the implementation of our method on the PV-RCNN backbone, with RP-PV (T2) as the teacher detector. Finally, referring to the experimental settings of RPKD-PV, we performed generalization verification of our method on the Voxel-RCNN backbone with RPKD-Voxel, using RP-Voxel (T3) as the teacher detector.
| Data Type | Detection Method | Vehicle(L1) | Vehicle(L2) | Pedestrian(L1) | Pedestrian(L2) | mAP/mAPH |
| Raw Data | PV-RCNN | 61.37/60.01 | 53.20/52.00 | 65.69/33.70 | 59.00/30.27 | 59.82/43.99 |
| RP-PV(T2) | 61.75/60.76 | 53.44/52.56 | 60.42/29.80 | 54.14/26.70 | 57.44/42.46 | |
| PCC-S-C12 (Bpp=4.637) | PV-RCNN(Baseline) | 61.27/60.07 | 52.87/51.81 | 59.89/30.98 | 53.60/27.78 | 56.91/42.66 |
| RPKD-PV(Ours) | 62.08/61.02 | 53.72/52.77 | 59.62/34.44 | 53.37/30.82 | 57.20/44.76 | |
| PCC-S-C11 (Bpp=2.711) | PV-RCNN(Baseline) | 60.51/59.52 | 52.30/51.43 | 58.75/29.72 | 52.57/26.61 | 56.03/41.82 |
| RPKD-PV(Ours) | 61.41/60.38 | 53.22/52.30 | 58.33/32.55 | 52.21/29.14 | 56.29/43.59 | |
| PCC-S-C10 (Bpp=1.359) | PV-RCNN(Baseline) | 60.09/58.81 | 51.95/50.82 | 56.76/26.30 | 50.77/23.53 | 54.89/39.86 |
| RPKD-PV(Ours) | 61.13/60.04 | 52.92/51.95 | 57.49/28.16 | 51.45/25.20 | 55.75/41.34 |
IV-B Experimental Results
Table I presents the performance improvement of our RPKD framework in 3D object detection for compressed point clouds with the PV-RCNN and Voxel-RCNN backbones on the KITTI validation split. For the compressed point clouds of PCC-S-C12, PCC-S-C11, and PCC-S-C10, RPKD-PV increases mAP by 3.33, 4.23, and 3.65, while RPKD-Voxel improves mAP by 2.25, 3.22, and 3.3. As a result, it reveals that the detectors of our RPKD framework exhibit consistent effectiveness and robustness across multiple code rates and backbones. Additionally, for compressed point clouds with the same code-rate order, RPKD-PV boosts the AP values for moderate-level pedestrians by 9.02, 11.98, and 9.66, respectively, and RPKD-Voxel elevates the AP values for moderate-level pedestrians by 7.93, 8.59, and 7.36. These findings highlight the significant enhancement in 3D detector robustness for challenging small-sized pedestrians within compressed point clouds.
| Layer Quantity | mAP | |||||
| 1 | 3 | |||||
| 61.31 | ||||||
| 61.81 | ||||||
| 62.59 | ||||||
| 63.70 | ||||||
As depicted in Fig. 5(a), based on the PV-RCNN backbone, we compare the detection performance of PV-RCNN (baseline), existing SparseKD, simulated RPV-KD and our RPKD on the KITTI compressed point clouds across multiple code rates. RPKD-PV achieves higher mAP values in both medium and high code rates, and its mAP value in the medium code rate exceeds that of RP-PV (T2). Fig. 5(b) shows the generalization performance of our RPKD framework on the Voxel-RCNN backbone. RPKD-Voxel outperforms Voxel-RCNN (baseline) in all code rates, and its mAP value surpasses that of RP-Voxel (T3) and Voxel-RCNN backbone in the high code rate. These results vividly illustrate the outstanding robustness and remarkable enhancement capabilities of our method in 3D object detection for compressed point clouds.
For the generalization ability in WOD, we evaluate the detection performance of our RPKD framework across multiple code rates with the PV-RCNN backbone on the validation split. In Table II, RPKD-PV improves mAP/mAPH for PCC-S-C12, PCC-S-C11, and PCC-S-C10 by 0.29/2.1, 0.26/1.77, and 0.86/1.48, respectively. For the compressed point clouds of PCC-S-C12, the APH values for vehicle and pedestrian categories exceed those of RP-PV (T2) and PV-RCNN backbone on raw data. The experimental results demonstrate that our RPKD framework maintains consistent effectiveness throughout diverse code rates, backbone networks and detection datasets.
| Options of RKD | Detectors of DKD | mAP | ||||
| Voxel-wise | Point-wise | PV-RCNN | RP-PV | |||
| 71.39 | ||||||
| 72.00 | ||||||
| 69.12 | ||||||
| 70.16 | ||||||
| 71.68 | ||||||
| 71.57 | ||||||
| 73.60 | ||||||
| RP Module | RKD | DKD | mAP |
| 69.37 | |||
| 72.00 | |||
| 71.68 | |||
| 73.60 |
IV-C Ablation Study
We conducted comprehensive ablation experiments with the PV-RCNN backbone to verify the effectiveness of our RPKD framework. In this section, we used raw point clouds from the KITTI dataset and their corresponding compressed point clouds of PCC-S-C11 as experimental data. The evaluation metric is the mAP of compressed point-cloud detectors on the KITTI validation split.
IV-C1 Different Configurations of RP Module
For non-reflectance compressed point-cloud voxels, the RP module employed 3D sparse convolution and voxel set extraction to generate point-wise geometric features for compressed point reflectance prediction. In terms of expanding receptive field, we modified the kernel size and padding of 3D sparse convolution, as well as the pooling radius of voxel set extraction. To assess the impact of network layer quantity on detection performance, we designed two network schemes: a single-layer network and a double-layer network . Additionally, we also adjusted the RP module’s loss weight for performance comparison. As shown in Table III, when using the single-layer network with and , the compressed point-cloud detector achieves optimal detection performance.
| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
| RPKD-PV(Unified) | 91.46 | 82.92 | 80.01 | 65.55 | 58.14 | 53.16 | 90.54 | 67.30 | 62.98 | 72.45 | |
| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
| RPKD-PV(Unified) | 91.20 | 82.37 | 79.29 | 66.71 | 59.32 | 54.18 | 87.66 | 66.86 | 62.26 | 72.21 | |
| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
| RPKD-PV(Unified) | 91.37 | 82.37 | 79.48 | 65.38 | 58.05 | 52.91 | 85.69 | 61.66 | 57.44 | 70.48 | |
IV-C2 Critical Elements of Our RPKD
As indicated in Table IV, we assessed the influences of critical elements in our RPKD framework on the detection performance for compressed point clouds. In the first two rows, we compared the pre-training performance of compressed point-cloud detectors, where refers to the voxel-based averaging operation in the RCM module. denotes screening foreground compressed points in calculating reflectance distillation loss. In the third row, we used the matched raw point reflectance as the distillation label during point-wise RKD and did not perform the voxel-based . In contrast, we utilized the matched reflectance of raw point-cloud voxels as the distillation label during voxel-wise RKD and conducted the voxel-based , as shown in the fourth row. In Table IV, PV-RCNN and RP-PV represent raw point-cloud detector types that provided the first-stage proposals for DKD. As indicated in the last row, our RPKD framework achieves the best performance with these settings.
IV-C3 Designed Modules of Our RPKD
As shown in Table V, we evaluated the designed module’s effectiveness in 3D object detection for compressed point clouds. The first row displays the detection performance of PV-RCNN (baseline) for compressed point clouds of PCC-S-C11. We then incrementally introduced the RP module, RKD and DKD into the detector and reported corresponding mAP values. All the designed modules improve the performance of PV-RCNN (baseline), with the best mAP value achieved when all modules are utilized together.
IV-C4 Experiments for a Unified Model
To develop a unified detector applicable to various code rates, we combined the KITTI training splits of PCC-S-C10, PCC-S-C11, and PCC-S-C12 in a ratio as input for the unified model. The trained model was then used to perform inference on each validation split of these compressed point clouds to evaluate detection performance. As shown in Table VI, RPKD-PV (unified) increases mAP by 2.84 for both PCC-S-C12 and PCC-S-C11, and by 4.04 for PCC-S-C10. These results demonstrate that our RPKD framework under a unified model achieves consistent effectiveness across diverse code rates.
IV-C5 Visual Comparisons of Detection Results
As depicted in Fig. 6, we performed visual comparisons of different data types and detection methods on the KITTI validation split. For compressed point-cloud detectors, RPKD-PV achieves fewer false positives and missed detections across multiple code rates, demonstrating the stronger robustness of our RPKD framework in 3D object detection for compressed point clouds. However, compared to the raw point-cloud detector based on PV-RCNN, RPKD-PV tends to miss objects with fewer points or at farther distances, which reveals a limitation of our method in detecting such challenging objects.
IV-C6 Visual Comparisons of Reflectance Prediction
In each BEV grid, we selected the maximum reflectance value from either raw data or compressed point clouds for coloring the corresponding visual points. As shown in Fig. 7, the predicted reflectance in compressed point clouds closely matches the actual reflectance in raw data within the major reflective areas of foreground objects. These results demonstrate that our RP module accurately predicts the reflectance of foreground objects across various code rates, thereby improving detection robustness for compressed point clouds.
IV-D Discussion of Limitation and Future Work
For non-reflectance compressed point clouds, this paper effectively supplements the missing reflectance information through the RP module and RKD, thus enhancing detection robustness. However, as shown in Table I and Table II, the APs for pedestrians in WOD and for moderate-level cars and bicycles in KITTI still fall short of the corresponding APs achieved with raw point clouds. This may be because RPKD focuses on transferring detection knowledge via DKD without performing explicit geometric enhancement on lossy compressed point clouds. In future work, we plan to explore geometric enhancement modules at the receiver to recover missing spatial details, thereby further improving the robustness of 3D object detection against non-reflectance compressed point clouds.
V Conclusion
We have presented a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds. During detection training, the designed RCM module assigns the average reflectance of raw point-cloud voxels to corresponding compressed points based on their spatial relationship, providing essential prior labels for reflectance reconstruction. Treating non-reflectance compressed point clouds as weak students, the geometry-based RP module plays a critical role in precise reflectance prediction and object detection within compressed point-cloud detectors. The raw point-cloud detector, with the same structure, transfers the accurate reflectance prediction and object detection knowledge through RKD and DKD, significantly enhancing the robustness of compressed point-cloud detectors. Experimental results on the KITTI dataset and WOD demonstrate that our RPKD framework consistently improves detection performance for compressed point clouds across various code rates and backbone networks.

### v2

@@
 Abstract
-Regarding intelligent transportation systems for vehicle networking, low-bitrate transmission via lossy point cloud compression is vital for facilitating real-time collaborative perception among vehicles with restricted bandwidth.
+Regarding intelligent transportation systems, low-bitrate transmission via lossy point cloud compression is vital for facilitating real-time collaborative perception among connected agents, such as vehicles and infrastructures, under restricted bandwidth.
 In existing compression transmission systems, the sender lossily compresses point coordinates and reflectance to generate a transmission code stream, which faces transmission burdens from reflectance encoding and limited detection robustness due to information loss.
@@
 The discarded reflectance is then reconstructed by a geometry-based reflectance prediction (RP) module within the student detector for precise detection.
-A teacher detector with the same structure as student detector is designed for performing reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) from raw to compressed point clouds.
-Our RPKD framework jointly trains detectors on both raw and compressed point clouds to improve the student detector’s robustness.
-Experimental results on the KITTI dataset and Waymo Open Dataset demonstrate that our method can boost detection accuracy for compressed point clouds across multiple code rates.
-Notably, at a low code rate of 2.146 Bpp on the KITTI dataset, our RPKD-PV achieves the highest mAP of 73.6, outperforming existing detection methods with the PV-RCNN baseline.
-Index Terms:
-Compressed Point Clouds, 3D Object Detection, Knowledge Distillation, Reflectance Prediction.I Introduction
+A teacher detector with the same structure as the student detector is designed for performing reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) from raw to compressed point clouds.
+Our cross-source distillation training strategy (CDTS) equips the student detector with robustness to low-quality compressed data while preserving the accuracy benefits of raw data through transferred distillation knowledge.
+Experimental results on the KITTI and DAIR-V2X-V datasets demonstrate that our method can boost detection accuracy for compressed point clouds across multiple code rates.
+We will release the code publicly at https://github.com/HaoJing-SX/RPKD.
+I Introduction
 LiDAR point clouds, composed of sparse points with both geometric (coordinates) and attribute (reflectance) information about object surfaces, are vital for predicting the categories, positions, and bounding boxes of 3D objects.
-In cooperative perception among networked vehicles, local vehicles (receivers) need to rapidly acquire LiDAR point clouds captured by neighboring vehicles (senders) to enable an expanded perception range and accurate object detection [1, 2, 3, 4].
+In vehicle-to-everything (V2X) cooperative perception, receivers need to rapidly acquire LiDAR point clouds captured by neighboring senders to enable an expanded perception range and accurate object detection [1, 2, 3, 4].
 Raw point clouds are extremely voluminous and require a high transmission bandwidth, making it difficult to realize real-time cooperative perception.
@@
 Therefore, enhancing the 3D detector’s robustness for compressed point clouds is of great research significance.
-Existing 3D object detection methods in compressed point clouds can be classified into two categories based on their reflectance processing approaches: 3D object detection with or without reflectance encoding.
-For 3D object detection with reflectance encoding, the geometry-based point cloud compression (G-PCC) [5] separately compresses the coordinates and reflectance of voxelized point clouds to form geometry and attribute bitstreams for reconstruction and detection tasks, as shown in Fig.
+Existing 3D object detection methods in compressed point clouds can be classified into two categories based on their reflectance processing approaches: 3D object detection with [5, 6, 7] or without reflectance encoding [8, 9, 10, 11, 12, 13].
+For 3D object detection with reflectance encoding, both G-PCC [5] and Unicorn [6, 7] compressed voxelized point-cloud coordinates together with reflectance, generating geometry and attribute bitstreams for reconstruction and detection tasks, as shown in Fig.
 1(a).
-Building on this, Unicorn [6, 7] introduces a learning-based compression framework that utilizes multiscale sparse tensors from single-frame voxelized point clouds and lower-scale priors from multi-frame point clouds to improve compression and reconstruction performance.
-However, these methods face a significant burden from the attribute bitstream during networking transmission, limiting transmission efficiency.
-To solve this issue, in 3D object detection without reflectance encoding, existing compression methods based on voxelized point clouds and deep learning [8, 9, 10, 11, 12, 13] typically focus on compressing geometric coordinates while discarding reflectance to elevate transmission efficiency, as presented in Fig.
+However, the attribute bitstream in these methods introduces a transmission burden that multiplies with the number of connected vehicles and infrastructures.
+To solve this issue, in 3D object detection without reflectance encoding, existing compression methods based on voxelized point clouds and deep learning [8, 9, 10, 11, 12, 13] typically focused on compressing geometric coordinates while discarding reflectance to elevate transmission efficiency, as presented in Fig.
 1(b).
-As an illustration, PCC-S [9] employs a geometric perception module to improve the learning of sibling and ancestor contexts in an octree structure for excellent geometric coordinates compression and reconstruction, during which the decoded data is termed non-reflectance compressed point clouds.
 However, lossy compression by these methods causes substantial information loss in low-bitrate transmission, severely impacting 3D object detection accuracy.
-To address these challenges, this paper aims to improve the transmission efficiency of lossy compressed point clouds and develop robust compressed point-cloud detectors, supported by a series of detailed studies.
-Fig.
-1(c) shows a baseline of our method.
-It takes accurate object predictions from raw point clouds as teacher knowledge for distillation during detection training.
-In this setup, the teacher model transfers its knowledge to compressed point-cloud detectors through response-based detection knowledge distillation (DKD).
-As shown in Fig.
-2, SparseKD-PV outperforms PV-RCNN by an average of 1.11% in mAP across multiple data types, whereas it still exhibits a significant detection performance gap compared to PV-RCNN’s mAP of 75.56 on raw data.
-In general, LiDAR point cloud reflectance provides crucial cues for identifying 3D objects, thus improving the reliability of 3D object detection.
-In SparseKD-PV, the decision to discard reflectance information to enhance transmission efficiency is a key factor contributing to the suboptimal performance.
-While methods like Unicorn [7] reconstruct reflectance via a residual-based attribute bitstream using a DNN network, and PCC-S [9] does not transmit the attribute bitstream but needs alternative strategies to recover reflectance, we argue that relying solely on non-reflectance compressed point clouds for detection misses important reflectance features.
-To address this limitation, we propose a dedicated reflectance prediction (RP) module at the receiver, which utilizes the geometric information of compressed point clouds and reflectance priors from raw point clouds.
-Additionally, to further enhance the accuracy of reflectance prediction, we introduce response-based reflectance knowledge distillation (RKD), which effectively guides the RP module during learning process, improving detection performance while maintaining efficient data transmission.
+To address these challenges, this paper adopts the voxel-based lossy geometric compression method PCC-S [9] to maximize bandwidth reduction and systematically investigates training strategies for non-reflectance compressed point clouds to develop a robust detector.
+The existing raw-data single-source training strategy (STS-R) [8, 14] pretrained detection models on raw data and evaluated them on low-quality compressed data, resulting in limited robustness.
+To handle the low quality of compressed point clouds, we adopt a compressed-data single-source training strategy (STS-C), training and evaluating at each compression rate to improve detection performance.
+In 3D object detection knowledge distillation, FD [15] transferred classification, regression, and feature knowledge from high-resolution to low-resolution data, thereby enhancing detection accuracy on low-resolution inputs.
+Building on this idea, we propose a cross-source distillation training strategy (CDTS), in which raw and compressed point-cloud models form a teacher-student model pair, and response-based detection knowledge distillation (DKD) from SparseKD [16] is applied to transfer accurate detection knowledge, as illustrated in Fig.
+1(c).
+Using PV-RCNN as the backbone, Fig.
+2 reports performance for different training strategies and detection methods on non-reflectance compressed data.
+SparseKD-PV (CDTS) outperforms PV-RCNN (STS-C) by an average of 1.11 mAP across various data types, yet still lags behind the PV-RCNN baseline on raw data.
+In general, LiDAR reflectance offers critical cues for 3D object recognition, thereby enhancing detection reliability.
+For SparseKD-PV (CDTS), however, discarding reflectance to improve transmission efficiency remains a key factor contributing to the suboptimal performance.
+To overcome this limitation, we introduce a dedicated reflectance prediction (RP) module at the receiver and employ knowledge distillation techniques to refine reflectance estimation, ultimately boosting detection performance on compressed point clouds.
 Incorporating the previous analysis, we propose a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds, as depicted in Fig.
@@
 Besides, the proposed reflectance prediction (RP) module comprehensively extracts geometric features of voxelized point clouds, achieving reflectance reconstruction at the receiver.
-Through our reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD), rich teacher knowledge is transferred from raw to compressed point clouds for more precise reflectance reconstruction and robust 3D object detection.
+Our cross-source distillation training strategy (CDTS) utilizes response-based reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) to transfer guided knowledge from raw to compressed point-cloud detectors, enabling more accurate reflectance reconstruction and improved 3D object detection.
 As shown in Fig.
-2, RPKD-PV significantly increases the mAP values across multiple data types, which demonstrates the effectiveness of our RPKD framework for non-reflectance compressed point clouds.
+2, RPKD-PV (CDTS) significantly raises the mAP values across various data types, demonstrating the effectiveness of our method for non-reflectance compressed point clouds.
 In summary, the contributions of this work are as follows.
@@
 At the receiver, we design a geometry-based RP module to reconstruct discarded reflectance, thereby improving detection performance for non-reflectance compressed point clouds.
-We introduce two response-based distillation constraints, namely RKD and DKD, to transfer supervisory knowledge from raw to compressed point-cloud detectors, further enhancing detection robustness for compressed point clouds.
+The proposed CDTS transfers supervisory knowledge from raw to compressed point-cloud detectors using two response-based distillation constraints, RKD and DKD, further enhancing detection robustness for low-quality compressed data.
 The remainder of the paper is organized as follows.
@@
 In Section III, we present the overall architecture of a 3D object detection framework with reflectance prediction-based knowledge distillation and elaborate on its key components.
-Section IV demonstrates the effectiveness of our method on the KITTI dataset and Waymo Open Dataset and includes ablation studies to evaluate the effect of each component.
+Section IV demonstrates the effectiveness of our method on the KITTI and DAIR-V2X-V datasets, and includes ablation studies to assess the impact of each component.
 Finally, Section V concludes this paper.
@@
 II-A 3D Object Detection
-In 3D object detection for raw point clouds, existing methods can be sorted into two classes according to the data format of network input: point-based [14, 15, 16, 17, 18, 19, 20, 21] and voxel-based [22, 23, 24, 25, 26, 27, 28, 29, 30, 31] methods.
+In 3D object detection for raw point clouds, existing methods can be sorted into two classes according to the data format of network input: point-based [17, 18, 19, 20, 21, 22] and voxel-based [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] methods.
 Point-based detectors extract hierarchical point-wise features by various sampling and aggregation methods to predict 3D objects.
 In contrast, voxel-based detectors utilize 3D sparse convolution to extract voxel features, and these features are projected onto the bird’s eye view (BEV) for 3D object detection.
-In 3D object detection for compressed point clouds, those generated by range image-based [32, 33, 34, 35] and voxel-based [8, 9] compression frameworks are input into compressed point-cloud detectors for recognition and localization.
-These methods generally rely on raw point clouds for pre-training 3D detectors and assess detection performance on lossy compressed point clouds across various code rates using the pre-trained models, as shown in Fig.
-1(a) and 1(b).
-However, they lack effective robustness measures against information loss and reconstruction errors caused by lossy point cloud compression, leading to a significant decline in detection accuracy.
-For robust 3D object detection in compressed point clouds, this paper addresses three key aspects.
-First, voxelized representation of compressed point clouds helps mitigate the negative effects of lossy compression.
-Thus, we develop robust compressed point-cloud detectors on voxel-based detection frameworks.
-Second, pre-training 3D detectors solely on raw point clouds is insufficient for adapting to the characteristics of compressed point clouds.
-To improve detection performance, our 3D detectors are trained on both raw and compressed point clouds.
-Finally, with a two-source joint training scheme, we use voxel-based two-stage detectors like PV-RCNN [22] and Voxel-RCNN [23] as backbone networks and transfer detection knowledge of the first-stage proposals from raw to compressed point-cloud detectors, thereby enhancing detection robustness for compressed point clouds.
+In 3D object detection for compressed point clouds, those generated by range image-based [35, 14, 36, 37] and voxel-based [5, 6, 7, 8, 9, 10, 11, 12, 13] compression frameworks were input into detectors for recognition and localization.
+These methods typically rely on raw point clouds for pretraining 3D detectors and evaluate detection performance on lossy compressed point clouds at various code rates using the pretrained models.
+However, the raw-data single-source training strategy lacks adaptability to low-quality compressed data, resulting in a significant decline in detection accuracy.
+To address this, we propose a novel CDTS that improves detector robustness to low-quality compressed data while inheriting the accuracy of raw data.
 II-B Point Cloud Compression
@@
 However, these methods are confronted with uneven 3D object loss and noise-point interference due to the lack of spatial structure and context information during lossy compression.
-Image-based methods project point clouds onto 2D images for quantization and encoding, but they introduce more severe three-dimensional spatial distortion and information loss compared to voxel-based methods.
+Image-based methods project point clouds onto 2D images for quantization and encoding, but they introduce more severe three-dimensional spatial distortion and reconstruction errors compared to voxel-based methods.
 By analyzing octree or multi-scale sparse tensors, voxel-based methods enhance compression performance via context feature learning, multi-scale joint optimization, and redundant encoding reduction.
-Among these compression methods, only G-PCC [5] and Unicorn [7] incorporate reflectance encoding modules, as reflectance encoding presents challenges in structural design, end-to-end optimization, and data-flow overhead.
+Among these compression methods, only G-PCC [5] and Unicorn [7] incorporated reflectance encoding modules, as reflectance encoding presents challenges in structural design, end-to-end optimization, and data-flow overhead.
 To solve this problem, we design an RP module at the receiver to supplement reflectance information for detection training on non-reflectance compressed point clouds generated by voxel-based methods.
@@
 Knowledge distillation methods for 3D object detection can be divided into three categories according to their research objectives: model compression distillation [16, 47, 48, 49, 50, 51], detection improvement distillation [52, 53], and weak-data robustness distillation [54, 15, 55].
-In model compression distillation, SparseKD [45] introduces response-based constraints using Logit and Label KD to guide lightweight student models, achieving accurate and efficient 3D object detection.
-In detection improvement distillation, SPNet [51] assigns category values to the interior points of ground truth (GT) boxes, while X-Ray [52] addresses object point loss caused by sparsity and occlusion utilizing tracking information from consecutive-frame point clouds.
+In model compression distillation, SparseKD [16] introduced response-based constraints through Logit and Label KD to guide lightweight student models, achieving accurate and efficient 3D object detection.
+In detection improvement distillation, SPNet [52] assigned category values to the interior points of ground truth (GT) boxes, while X-Ray [53] addressed object point loss caused by sparsity and occlusion utilizing tracking information from consecutive-frame point clouds.
 Both methods augment data representations with strong semantic priors, providing valuable supervisory knowledge for distillation.
-In weak-data robustness distillation, student models employ low-quality point clouds from rainy conditions [55] or low-resolution sensor data [53, 54], learning complete and accurate knowledge from raw point clouds to improve detection robustness in weak data scenarios.
-Inspired by these weak-data robustness distillation methods, our RPKD framework treats non-reflectance compressed point clouds involved in vehicle networking data transmission as weak data for student models and raw point clouds as guiding data for teacher models, aiming to develop robust compressed point-cloud detectors.
+In weak-data robustness distillation, student models employed low-quality point clouds from rainy conditions [55] or low-resolution sensor data [54, 15], learning complete and accurate knowledge from raw point clouds to improve detection robustness in weak-data scenarios.
+Inspired by weak-data robustness distillation methods, our RPKD framework treats non-reflectance compressed point clouds in vehicle networking data transmission as weak data for student models, with raw point clouds guiding teacher models, aiming to develop robust compressed point-cloud detectors.
+II-D V2X Collaborative Perception
+To overcome single-vehicle perception limitations, V2X collaborative perception [3, 4, 56, 57, 58] develops early, intermediate, and late fusion strategies based on shared information among connected agents.
+These strategies correspond to the transmission of raw data, intermediate layer features, and prediction results, respectively.
+In early fusion with LiDAR [1, 2], raw point clouds offer precise scene information and are easily registered, but they require significant transmission bandwidth.
+Cooper [1] mitigated the bandwidth demand for mutual transmission between two vehicles by extracting key point clouds using RoIs.
+However, this approach still faces challenges in scaling to large V2X collaborative scenarios.
+To address these issues, this paper pioneeringly introduces voxel-based lossy geometric compression into collaborative perception, effectively balancing low bandwidth and high accuracy.
+Regarding datasets, DAIR-V2X [58] is more mature and was released earlier than V2X-Radar [56].
+For detection tasks, we focus on single-view 3D detection in DAIR-V2X-V and evaluate performance on compressed data across various object categories, thus laying the foundation for future collaborative detection.
 III Proposed Approach
 III-A Overview of our Framework
-During point cloud compression, reflectance encoding faces significant challenges in both end-to-end optimization and data-flow overhead.
+In collaborative perception with LiDAR point cloud compression, reflectance encoding imposes a linearly increasing transmission burden as the number of connected agents grows.
 According to the separable compression mechanism of G-PCC [5], existing methods typically disregard reflectance compression and focus solely on optimizing geometric coordinate encoding to achieve efficient transmission and accurate detection.
@@
 First, our reflectance cross-match (RCM) module correlates lossy compressed points with neighboring raw point-cloud voxels.
-The nearest voxel’s average reflectance is assigned as the reflectance label for the current compressed point.
-Similarly, the reflectance inter-match (RIM) module assigns a voxel’s average reflectance as the reflectance label for corresponding raw points.
+Average reflectance of the nearest voxel is assigned as the reflectance label for the current compressed point.
+Similarly, the reflectance inter-match (RIM) module assigns the average reflectance of a voxel as the label for corresponding raw points.
 Next, the proposed reflectance prediction (RP) module generates point-wise geometric features using 3D sparse convolution and voxel set extraction to predict reflectance.
-The predicted point-wise reflectance is then incorporated into corresponding non-reflectance voxels for subsequent detection tasks.
-Finally, we treat non-reflectance compressed point clouds as weak data and introduce two response-based knowledge distillation constraints: reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD).
-Specifically, RKD transfers reflectance prediction knowledge from raw point clouds via the RP module, while DKD conveys first-stage proposal knowledge from the two-stage 3D detector for raw point clouds.
+The predicted reflectance is then integrated into the corresponding non-reflectance voxels for subsequent detection tasks.
+Finally, our cross-source distillation training strategy (CDTS) transfers guiding knowledge from raw to compressed point-cloud detectors through two response-based distillation constraints.
+Specifically, reflectance knowledge distillation (RKD) and detection knowledge distillation (DKD) concentrate on the RP module’s reflectance knowledge and the first-stage proposal knowledge, respectively.
 In what follows, we detail the proposed framework.
 III-B Reflectance Label Generation
-Due to point number reduction and reconstruction errors from lossy compression, compressed points do not correspond one-to-one with raw points.
+Due to point number reduction and reconstruction errors from voxel-based lossy compression, compressed points do not correspond one-to-one with raw points.
 For compressed point reflectance prediction, we introduce an RCM module based on spatial relationships to assign reflectance labels from raw point-cloud voxels to corresponding compressed points.
@@
 For geometric coordinate encoding, voxel-based compression methods [6, 9] are often employed to obtain non-reflectance compressed point clouds.
-Before training, with each point’s reflectance assigned to 0, the compressed point clouds fed into our detection networks are .
-Raw point clouds are represented as .
-These point clouds are voxelized to the same scale, generating compressed point-cloud voxels and raw point-cloud voxels .
-When using , and as inputs, the principle of our RCM module can be described as follows.
+With each compressed point’s reflectance assigned to zero, raw and compressed point clouds fed into our detection networks are represented as and .
+These point clouds are voxelized to the same scale, generating raw point-cloud voxels and compressed point-cloud voxels .
+When using the compressed point coordinates and the voxels of raw and compressed point clouds as inputs, the principle of our RCM module can be described as follows.
 As depicted in Fig.
 4, there are two spatial matching strategies between compressed points and raw point-cloud voxels: center matching and surrounding matching.
-Specifically, the voxel corresponding to compressed point coordinates is denoted as , and the surrounding non-empty voxels within a 1-voxel matching range are .
+Specifically, the center voxel corresponding to compressed point coordinates is denoted as , and the surrounding non-empty voxels within a 1-voxel matching range are .
 The initial reflectance labels for compressed points are defined as:
-where represents the average reflectance of in center matching, and is the nearest voxel’s average reflectance in surrounding matching.
-However, when multiple compressed points exist within the same compressed point-cloud voxel , their reflectance labels may come from different raw point-cloud voxels, resulting in significant numerical discrepancies that negatively impact compressed point reflectance prediction.
+where represents the average reflectance of a center voxel, and is the nearest voxel’s average reflectance in surrounding matching.
+However, when multiple compressed points exist within a compressed point-cloud voxel, their reflectance labels may come from different raw point-cloud voxels, resulting in significant numerical discrepancies that negatively impact compressed point reflectance prediction.
 To address this issue, we further calculate the average reflectance labels for all points within each compressed point-cloud voxel.
 The final reflectance labels for compressed points are thus expressed as:
-where denotes the reflectance label for each point within the current compressed point-cloud voxel , and represents the averaging operation applied to these reflectance labels.
+where denotes the reflectance label for each point within the current compressed point-cloud voxel , and represents the averaging operation applied to these reflectance labels, with and denoted as the corresponding numbers of compressed points and voxels, respectively.
 Reflectance Inner-match Module.
-In the RPKD framework, it is crucial to establish suitable reflectance prediction knowledge from raw point clouds to guide the prediction process for compressed point-cloud reflectance.
-Specifically, we treat the reflectance predictions for compressed and raw point clouds as a student-teacher pair, where both sets of labels need to maintain spatial correspondence and information similarity.
+In the RPKD framework, it is crucial to establish suitable reflectance knowledge from raw point clouds to guide the prediction process for compressed point-cloud reflectance.
+Specifically, we treat the reflectance predictions for raw and compressed point clouds as a teacher-student pair, where both sets of labels need to maintain spatial correspondence and information similarity.
 In the RCM module, the reflectance labels for compressed points are derived from the corresponding raw point-cloud voxels.
-Therefore, our RIM module avoids directly using raw point reflectance and instead assigns the average reflectance of raw point-cloud voxels to corresponding points as their reflectance labels.
-Since raw points exactly correspond to their voxels, our RIM module employs center matching.
-With raw point coordinates and voxels as inputs, the raw point reflectance labels are computed by the following formula:
-where is the voxel corresponding to raw point coordinates , and denotes the average reflectance of this voxel.
+Therefore, our RIM module avoids directly using raw point reflectance and instead assigns the average reflectance of raw point-cloud voxels to corresponding points.
+Since raw points correspond exactly to their voxels, the RIM module uses center matching to obtain the raw point reflectance labels .
+These pluggable matching modules are designed for non-reflectance compressed point clouds, which are obtained through voxel-based geometric compression methods such as PCC-S.
+When compressed point clouds inherently include reflectance, the reflectance labels are directly available.
 III-C Reflectance Prediction Module
-Concerning the reflectance reconstruction of compressed point clouds, Unicorn [7] utilizes current-scale reflectance residual features and low-scale reflectance priors through a DNN module to predict reflectance.
-In contrast, our RP module exploits the geometric features of non-reflectance compressed point clouds and the reflectance priors of raw point clouds to achieve precise reflectance reconstruction at the receiver.
-This innovative design not only reduces the transmission rate in compression transmission systems but also enhances the robustness of 3D detectors against non-reflectance compressed point clouds.
+Concerning the reflectance reconstruction of compressed point clouds, Unicorn [7] utilized current-scale reflectance residual features and low-scale reflectance priors through a DNN module to predict reflectance.
+In contrast, our RP module exploits the geometric features of non-reflectance compressed point clouds and the reflectance priors of raw point clouds to reconstruct reflectance at the receiver.
+In collaborative perception with LiDAR point cloud compression, the concept of reflectance prediction is crucial for ensuring detection robustness against low-quality compressed data, while eliminating reflectance encoding to push the boundaries of transmission bandwidth reduction.
 As shown in Fig.
@@
 The voxel set extraction structure in our RP module is identical to that in PV-RCNN [23], but the number of sampling points is equal to the number of compressed points in the current frame.
-Since the reflectance is a single-channel floating-point scalar in the range of , we use two fully connected (FC) layers as a prediction head to generate compressed point reflectance, denoted as .
-The predicted reflectance of compressed points, , is then incorporated into the corresponding compressed point-cloud voxels based on their spatial relationship for subsequent training and detection.
+Since the reflectance input to the detection network is a single-channel floating-point scalar in the range of , we use two fully connected (FC) layers as a prediction head to generate compressed point reflectance.
+The predicted reflectance is then incorporated into the corresponding compressed point-cloud voxels based on their spatial relationship for subsequent training and detection.
 We screen the compressed point reflectance labels with values greater than zero, and define their indices as .
-We apply Mean Squared Error Loss (MSELoss) to constrain compressed point reflectance prediction, and the corresponding loss function is expressed as follows:
-where represents the number of compressed point reflectance labels with values greater than zero.
+Mean Squared Error Loss (MSELoss) is applied to constrain compressed point reflectance prediction, and the corresponding loss function is expressed as follows:
+where represents the number of compressed point reflectance labels with values greater than zero, and denotes the predicted reflectance for compressed points.
 RP Module for Raw Point Clouds.
@@
 Non-reflectance raw point-cloud voxels are fed into this module, and the geometric features of raw points are represented as:
-Subsequently, we predict raw point reflectance and incorporate it into corresponding voxels to obtain raw point-cloud voxels with the recovered reflectance .
+where represents raw point coordinates, and is the number of raw points.
+Subsequently, we predict raw point reflectance and assign it to corresponding voxels, resulting in raw point-cloud voxels with recovered reflectance.
 We screen the raw point reflectance labels that are greater than zero and define their indices as . The loss function for raw point reflectance prediction is defined as:
-where represents the number of raw point reflectance labels with values greater than zero.
-III-D Reflectance Prediction-based Knowledge Distillation
-In lossy compression, non-reflectance compressed point clouds undergo point number reduction and reconstruction errors, leading to incomplete and biased geometric information.
-If only these compressed point clouds are used for reflectance prediction and object detection, it is difficult to match the detection accuracy of raw point clouds.
-Drawing from the weak-data robustness research [53, 54, 55] in 3D object detection knowledge distillation, we treat the non-reflectance compressed point clouds as weak students, enabling them to learn accurate and complete knowledge from raw point clouds to improve detection performance for compressed point clouds.
-Specifically, we propose two components: RKD, which transfers reflectance prediction knowledge from our RP module, and DKD, which refines the first-stage proposals in the two-stage 3D object detector, as depicted in Fig.
-3.
-Both proposed distillation methods fall under the response-based knowledge distillation category, and their principles are outlined as follows.
+where represents the number of raw point reflectance labels with values greater than zero, and denotes the predicted reflectance for raw points.
+III-D Reflectance Prediction-based CDTS
+Inspired by weak-data robustness research [54, 15, 55], the proposed CDTS employs raw and compressed data models in a teacher–student framework for knowledge distillation (as shown in Fig.
+5), thereby improving detection performance on low-quality compressed data.
+To tackle the critical challenge of missing LiDAR reflectance in lossy geometric compression, our RPKD (CDTS) strengthens reflectance reconstruction through the integration of reflectance priors, the RP module, and the RKD constraint.
+In knowledge distillation, we introduce two components: RKD, which transfers reflectance prediction knowledge from the RP module, and DKD, which refines first-stage proposals in the two-stage 3D object detector.
+Both methods belong to response-based knowledge distillation, and their principles are outlined below.
 Reflectance Knowledge Distillation.
-Building upon the RP module’s prediction outputs, we obtain the compressed point reflectance and raw point reflectance .
+Building upon the RP module’s outputs, we obtain the predicted raw and compressed point reflectance.
 Since they do not have a one-to-one correspondence, we map the recovered teacher reflectance from raw point-cloud voxels to corresponding compressed points.
 The mapped teacher reflectance of each compressed point serves as a soft label of RKD.
-According to the spatial relationship in the RCM module, the raw point-cloud voxel corresponding to compressed point coordinates is denoted as , and the RKD soft labels are obtained by the following formula:
-where represents the mapped reflectance value from a raw point-cloud voxel to corresponding compressed point coordinates , and is the number of compressed points that can be successfully matched with raw point-cloud voxels in the RCM module.
-We also adopt MSELoss as the loss function to constrain the reflectance distillation process, and the RKD loss function is formulated as
-where denotes screening foreground compression points during loss computation, and is the number of foreground compression points.
+We also adopt MSELoss to constrain the reflectance distillation process, and the RKD loss function is formulated as
+where denotes screening foreground compression points during loss computation, is the number of compressed points that can be successfully matched with raw point-cloud voxels, and is the number of foreground compression points.
 Detection Knowledge Distillation.
@@
 Additionally, the first-stage dense proposals passed to the second stage inherently carry valuable detection knowledge for distillation.
-DKD combines the compressed point-cloud proposals and raw point-cloud proposals to form a student-teacher pair for response-based knowledge distillation.
-In these proposals, and represent the sigmoid-activated classification scores and bounding boxes, respectively.
+DKD combines the raw and compressed point-cloud proposals to form a teacher-student pair for response-based knowledge distillation.
 For distillation constraints, we apply Logit KD proposed in SparseKD [16] to design the distillation loss.
@@
 III-E Loss Function
-In this study, we pre-train separate detectors for the compressed and raw point clouds and initialize the training parameters of our RPKD framework using these pre-trained models.
-Therefore, the compressed point-cloud detection loss and the raw point-cloud detection loss are defined as follows:
+In this study, we pretrain separate detectors on compressed and raw point clouds and initialize the RPKD framework parameters with these pretrained models. Their loss functions are defined as follows:
 where , , and represent the reflectance prediction, classification, bounding box and orientation losses of the first-stage detector, respectively.
@@
 IV Experiments
-For 3D object detection in compressed point clouds, we utilized LiDAR point clouds from the KITTI dataset and Waymo Open Dataset (WOD) as raw point-cloud inputs for the compression network.
-Subsequently, we compressed and reconstructed the geometric coordinates of these point clouds at multiple code rates through deep learning-based PCC-S, resulting in non-reflectance compressed point clouds.
-The proposed RPKD framework was trained on both compressed and raw point clouds, using only compressed point clouds for inference.
-Extensive experiments demonstrate that our method significantly enhances detection robustness for non-reflectance compressed point clouds generated by voxel-based compression.
-The following sections provide a comprehensive explanation of the implementation details, result comparisons, and ablation analysis.
-| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
-| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
-| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
-| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
-| Voxel-RCNN | 92.55 | 84.65 | 82.41 | 66.07 | 59.30 | 53.73 | 92.81 | 73.55 | 68.97 | 74.89 | |
-| RP-Voxel(T3) | 92.18 | 84.12 | 81.75 | 62.04 | 56.11 | 51.05 | 93.67 | 72.87 | 68.41 | 73.58 | |
-| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
-| RPKD-PV(Ours) | 92.02 | 83.13 | 80.21 | 65.72 | 58.68 | 53.72 | 90.39 | 68.63 | 63.92 | 72.94 | |
-| Voxel-RCNN(Baseline) | 92.27 | 82.79 | 80.17 | 63.64 | 56.21 | 51.21 | 90.69 | 71.35 | 66.92 | 72.81 | |
-| RPKD-Voxel(Ours) | 92.06 | 82.91 | 79.98 | 70.40 | 64.14 | 57.63 | 92.26 | 70.31 | 65.84 | 75.06 | |
-| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
-| RPKD-PV(Ours) | 91.38 | 82.95 | 79.96 | 70.93 | 62.81 | 56.39 | 89.21 | 66.68 | 62.09 | 73.60 | |
-| Voxel-RCNN(Baseline) | 92.03 | 82.72 | 80.16 | 61.49 | 54.77 | 49.30 | 88.27 | 67.99 | 63.97 | 71.19 | |
-| RPKD-Voxel(Ours) | 91.78 | 83.00 | 80.30 | 70.33 | 63.36 | 57.17 | 92.12 | 68.10 | 63.54 | 74.41 | |
-| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
-| RPKD-PV(Ours) | 91.53 | 82.52 | 79.45 | 63.24 | 57.24 | 51.90 | 85.51 | 61.70 | 57.68 | 70.09 | |
-| Voxel-RCNN(Baseline) | 90.96 | 81.37 | 78.73 | 60.10 | 51.66 | 45.88 | 83.05 | 62.86 | 58.62 | 68.14 | |
-| RPKD-Voxel(Ours) | 90.67 | 81.87 | 78.95 | 65.95 | 59.02 | 53.54 | 87.79 | 64.80 | 60.35 | 71.44 | |
+For 3D object detection in compressed point clouds, we employed LiDAR data from the KITTI and DAIR-V2X-V datasets as raw point-cloud inputs for the compression network.
+The geometric coordinates of these point clouds were compressed and reconstructed at multiple code rates using the voxel-based lossy geometric compression method PCC-S [9].
+Building on the proposed CDTS, extensive experiments demonstrate that our RPKD (CDTS) significantly enhances detection robustness for non-reflectance compressed point clouds.
+The following sections detail the implementation, comparative results, and ablation analysis.
+| Data Type | Bpp | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
+| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | ||||
+| Raw Data | - | PV-RCNN [23]* | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.55 |
+| PCC-S-C12 | 3.86 | PV-RCNN (STS-C) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
+| RPKD-PV (CDTS) | 92.02 | 83.13 | 80.21 | 65.72 | 58.68 | 53.72 | 90.39 | 68.63 | 63.92 | 72.94 | ||
+| PCC-S-C11 | 2.15 | PV-RCNN (STS-C) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
+| RPKD-PV (CDTS) | 91.38 | 82.95 | 79.96 | 70.93 | 62.81 | 56.39 | 89.21 | 66.68 | 62.09 | 73.60 | ||
+| PCC-S-C10 | 1.05 | PV-RCNN (STS-C) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
+| RPKD-PV (CDTS) | 91.53 | 82.52 | 79.45 | 63.24 | 57.24 | 51.90 | 85.51 | 61.70 | 57.68 | 70.09 | ||
+| Raw Data | - | Voxel-RCNN [24]* | 92.55 | 84.65 | 82.41 | 66.07 | 59.30 | 53.73 | 92.81 | 73.55 | 68.97 | 74.89 |
+| PCC-S-C12 | 3.86 | Voxel-RCNN (STS-C) | 92.27 | 82.79 | 80.17 | 63.64 | 56.21 | 51.21 | 90.69 | 71.35 | 66.92 | 72.81 |
+| RPKD-Voxel (CDTS) | 91.89 | 82.94 | 80.03 | 71.49 | 64.33 | 58.50 | 92.28 | 71.47 | 66.85 | 75.53 | ||
+| PCC-S-C11 | 2.15 | Voxel-RCNN (STS-C) | 92.03 | 82.72 | 80.16 | 61.49 | 54.77 | 49.30 | 88.27 | 67.99 | 63.97 | 71.19 |
+| RPKD-Voxel (CDTS) | 91.74 | 82.90 | 80.22 | 69.65 | 62.83 | 56.10 | 91.80 | 68.36 | 63.94 | 74.17 | ||
+| PCC-S-C10 | 1.05 | Voxel-RCNN (STS-C) | 90.96 | 81.37 | 78.73 | 60.10 | 51.66 | 45.88 | 83.05 | 62.86 | 58.62 | 68.14 |
+| RPKD-Voxel (CDTS) | 91.37 | 82.33 | 79.50 | 66.72 | 59.53 | 53.88 | 84.68 | 63.03 | 58.93 | 71.11 | ||
+| Raw Data | - | SECOND [25]* | 90.43 | 81.50 | 78.50 | 55.89 | 50.97 | 46.38 | 81.91 | 64.40 | 60.53 | 67.83 |
+| PCC-S-C12 | 3.86 | SECOND (STS-C) | 90.34 | 81.21 | 78.32 | 55.59 | 49.67 | 45.32 | 80.19 | 65.19 | 61.23 | 67.45 |
+| RPKD-SECOND (CDTS) | 91.08 | 81.96 | 78.87 | 58.42 | 53.06 | 48.55 | 86.00 | 68.38 | 64.35 | 70.07 | ||
+| PCC-S-C11 | 2.15 | SECOND (STS-C) | 90.17 | 80.90 | 78.09 | 52.58 | 47.66 | 43.16 | 81.92 | 61.21 | 57.31 | 65.89 |
+| RPKD-SECOND (CDTS) | 90.87 | 81.69 | 78.62 | 57.35 | 52.88 | 48.44 | 84.84 | 66.07 | 61.79 | 69.17 | ||
+| PCC-S-C10 | 1.05 | SECOND (STS-C) | 88.25 | 78.80 | 75.72 | 50.48 | 45.29 | 41.37 | 75.59 | 56.28 | 52.66 | 62.72 |
+| RPKD-SECOND (CDTS) | 90.08 | 80.74 | 77.52 | 56.76 | 52.34 | 47.71 | 81.81 | 62.39 | 58.32 | 67.52 | ||
 IV-A Implementation Details
 IV-A1 Datasets and Hardware
-For the KITTI dataset, we used 7481 labeled training samples as raw point clouds, dividing them into a training split with 3712 samples and a validation split with 3769 samples.
-Firstly, the PCC-S network employed the training split of raw point clouds for compression learning and reconstructed 7481 frames of point clouds through the trained compression model at octree levels of 12, 11, and 10, respectively.
-Regarding the training configuration, we trained the PCC-S network for 20 epochs on a single RTX 8000 GPU with a batch size of 2 and a maximum octree level of 12.
-Then, we established independent models for each code rate and conducted joint training on both the compressed and raw point clouds to build robust 3D detectors for compressed point clouds.
-These detectors were trained for 80 epochs on 4 RTX 4090 GPUs with a batch size of 16.
-WOD is a large-scale detection dataset composed of 798 training sequences and 202 validation sequences.
-Due to the limited training efficiency of PCC-S and the computing power of a single RTX 8000 GPU, we only selected 20 training sequences and 10 validation sequences for research.
-Thus, the raw point clouds consisted of a training split with 3966 samples and a validation split with 1967 samples.
-Using the same training configuration as KITTI, we obtained 5933 frames of compressed point clouds by PCC-S on WOD at octree levels 12, 11, and 10, respectively.
-Additionally, the detectors of our RPKD framework were trained for 30 epochs on 4 RTX 4090 GPUs with a batch size of 8.
+In KITTI, we selected 7,481 labeled training samples as raw point clouds, dividing them into a training split of 3,712 samples and a validation split of 3,769 samples.
+Similarly, in DAIR-V2X-V, we selected 6,509 training frames, which were split into 4,335 frames for training and 2,174 frames for validation.
+We trained the PCC-S network on the raw point-cloud training split for compression learning and applied to reconstruct point clouds at octree levels 12, 11, and 10, denoted as PCC-S-C12, PCC-S-C11, and PCC-S-C10, respectively.
+The compression network was optimized for 20 epochs on a single RTX 8000 GPU with a batch size of 2 and a maximum octree level of 12.
+Independent detection models for each code rate were trained for 80 epochs on 4 RTX 4090 GPUs with a batch size of 16.
 IV-A2 Training Parameters
 In the RCM module, we matched compressed points with raw point-cloud voxels within a 1-voxel matching range, corresponding to spatial voxels.
-During pre-training, the compressed and raw point-cloud detectors shared the same loss weights, and these loss weights were set to .
-In the knowledge distillation stage, we designed two distillation methods, RKD and DKD, whose loss weights were set to .
-Specifically, the distillation constraints for DKD consisted of classification distillation loss and bounding box distillation loss , and their loss weights were .
-For dataset configuration, we trained compressed point-cloud detectors using the standard parameters of public datasets and expanded single-source data to two-source data for joint processing.
+During pretraining, the raw and compressed point-cloud detectors shared the same loss weights, and these loss weights were set to .
+For CDTS, we introduced two distillation methods, namely RKD and DKD, whose loss weights were configured as .
+Specifically, the distillation constraints for DKD included classification distillation loss and bounding box distillation loss , and their loss weights were .
 IV-A3 Evaluation Metrics
-We evaluated the robustness of our RPKD framework for 3D object detection on compressed point clouds using the KITTI and WOD validation splits.
-The primary metric for KITTI was average precision (AP) at 40 recall positions with 3D IoU thresholds of 0.7, 0.5, and 0.5 for cars, pedestrians, and cyclists, respectively.
-For WOD, the metrics were AP and average precision weighted by heading (APH).
-To compare performance across multiple code rates and methods, we used mean Average Precision (mAP) and mean APH (mAPH) across categories and difficulty levels.
+We evaluated the robustness of our RPKD framework for 3D object detection on compressed point clouds using the KITTI and DAIR-V2X-V validation splits.
+For KITTI and DAIR-V2X-V, the primary metric is average precision (AP) at 40 recall positions with 3D IoU thresholds of 0.7, 0.5, and 0.5 for car, pedestrian, and cyclist categories, respectively.
 The Bit-per-point (Bpp) metric represents the average number of bits required to encode each point, indicating the code rate of compressed point clouds.
-IV-A4 Controlled Experiment Design
-Firstly, in terms of data types, raw data, PCC-S-C12, PCC-S-C11, and PCC-S-C10 represented raw point clouds, the compressed point clouds of octree levels 12, 11, and 10, respectively.
-Then, the detection performance of PV-RCNN at each code rate served as the baseline, and the designed control experiments involved three detection methods, namely SparseKD-PV, RPV-KD, and RPKD-PV.
-In SparseKD-PV, with the teacher detector of PV-RCNN (T1) on raw data, we utilized Logit KD and Label KD for distillation learning and performance evaluation of compressed point-cloud detectors.
-RPV-KD and SparseKD-PV share the same teacher model and distillation methods, while the key difference is that RPV-KD mapped raw point reflectance into corresponding compressed points based on the matching principle of our RCM module (Section III-B) to simulate compressed point clouds with decoded reflectance.
-RPKD-PV denotes the implementation of our method on the PV-RCNN backbone, with RP-PV (T2) as the teacher detector.
-Finally, referring to the experimental settings of RPKD-PV, we performed generalization verification of our method on the Voxel-RCNN backbone with RPKD-Voxel, using RP-Voxel (T3) as the teacher detector.
-| Data Type | Detection Method | Vehicle(L1) | Vehicle(L2) | Pedestrian(L1) | Pedestrian(L2) | mAP/mAPH |
-| Raw Data | PV-RCNN | 61.37/60.01 | 53.20/52.00 | 65.69/33.70 | 59.00/30.27 | 59.82/43.99 |
-| RP-PV(T2) | 61.75/60.76 | 53.44/52.56 | 60.42/29.80 | 54.14/26.70 | 57.44/42.46 | |
-| PCC-S-C12 (Bpp=4.637) | PV-RCNN(Baseline) | 61.27/60.07 | 52.87/51.81 | 59.89/30.98 | 53.60/27.78 | 56.91/42.66 |
-| RPKD-PV(Ours) | 62.08/61.02 | 53.72/52.77 | 59.62/34.44 | 53.37/30.82 | 57.20/44.76 | |
-| PCC-S-C11 (Bpp=2.711) | PV-RCNN(Baseline) | 60.51/59.52 | 52.30/51.43 | 58.75/29.72 | 52.57/26.61 | 56.03/41.82 |
-| RPKD-PV(Ours) | 61.41/60.38 | 53.22/52.30 | 58.33/32.55 | 52.21/29.14 | 56.29/43.59 | |
-| PCC-S-C10 (Bpp=1.359) | PV-RCNN(Baseline) | 60.09/58.81 | 51.95/50.82 | 56.76/26.30 | 50.77/23.53 | 54.89/39.86 |
-| RPKD-PV(Ours) | 61.13/60.04 | 52.92/51.95 | 57.49/28.16 | 51.45/25.20 | 55.75/41.34 |
+| Data Type | Bpp | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
+| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | ||||
+| Raw Data | - | PV-RCNN [23]* | 72.06 | 60.50 | 57.58 | 42.67 | 41.94 | 41.93 | 41.12 | 38.64 | 39.22 | 48.41 |
+| PCC-S-C12 | 4.01 | PV-RCNN (STS-C) | 71.76 | 60.15 | 57.22 | 37.65 | 36.14 | 36.09 | 34.57 | 34.24 | 34.65 | 44.72 |
+| RPKD-PV (CDTS) | 72.79 | 61.08 | 58.16 | 41.97 | 41.51 | 40.77 | 45.53 | 41.53 | 42.10 | 49.49 | ||
+| PCC-S-C11 | 2.13 | PV-RCNN (STS-C) | 70.49 | 58.79 | 55.48 | 30.79 | 30.98 | 30.85 | 34.00 | 32.09 | 32.42 | 41.77 |
+| RPKD-PV (CDTS) | 70.73 | 58.96 | 56.03 | 36.42 | 35.58 | 34.59 | 40.06 | 36.60 | 36.99 | 45.11 | ||
+| PCC-S-C10 | 0.96 | PV-RCNN (STS-C) | 68.13 | 56.49 | 52.06 | 8.52 | 8.77 | 8.96 | 12.86 | 13.18 | 12.43 | 26.82 |
+| RPKD-PV (CDTS) | 68.03 | 57.11 | 52.69 | 10.01 | 10.35 | 10.29 | 14.98 | 14.88 | 14.54 | 28.10 | ||
 IV-B Experimental Results
-Table I presents the performance improvement of our RPKD framework in 3D object detection for compressed point clouds with the PV-RCNN and Voxel-RCNN backbones on the KITTI validation split.
-For the compressed point clouds of PCC-S-C12, PCC-S-C11, and PCC-S-C10, RPKD-PV increases mAP by 3.33, 4.23, and 3.65, while RPKD-Voxel improves mAP by 2.25, 3.22, and 3.3.
-As a result, it reveals that the detectors of our RPKD framework exhibit consistent effectiveness and robustness across multiple code rates and backbones.
-Additionally, for compressed point clouds with the same code-rate order, RPKD-PV boosts the AP values for moderate-level pedestrians by 9.02, 11.98, and 9.66, respectively, and RPKD-Voxel elevates the AP values for moderate-level pedestrians by 7.93, 8.59, and 7.36.
-These findings highlight the significant enhancement in 3D detector robustness for challenging small-sized pedestrians within compressed point clouds.
+Table I presents the performance improvement of our RPKD framework on compressed point clouds, using the PV-RCNN, Voxel-RCNN, and SECOND backbones on the KITTI validation split.
+For PCC-S-C12, PCC-S-C11, and PCC-S-C10 compressed point clouds, RPKD-PV increases mAP by 3.33, 4.23, and 3.65, respectively; RPKD-Voxel improves mAP by 2.72, 2.98, and 2.97; and RPKD-SECOND achieves mAP gains of 2.62, 3.28, and 4.8.
+As a result, it reveals that our RPKD framework’s detectors exhibit consistent effectiveness and robustness across various code rates and backbones.
+Additionally, for compressed point clouds with the same code-rate order, RPKD-PV increases the AP values for moderate-level pedestrians by 9.02, 11.98, and 9.66, respectively; RPKD-Voxel boosts the AP values by 8.12, 8.06, and 7.87; and RPKD-SECOND improves the AP values by 3.39, 5.22, and 7.05.
+These findings highlight the remarkable enhancement in 3D detector robustness for challenging small-sized pedestrians in compressed point clouds.
 | Layer Quantity | mAP | |||||
@@
 As depicted in Fig.
-5(a), based on the PV-RCNN backbone, we compare the detection performance of PV-RCNN (baseline), existing SparseKD, simulated RPV-KD and our RPKD on the KITTI compressed point clouds across multiple code rates.
-RPKD-PV achieves higher mAP values in both medium and high code rates, and its mAP value in the medium code rate exceeds that of RP-PV (T2).
+6(a), we compared the detection performance of PV-RCNN (STS-C), SparseKD-PV (CDTS), and RPKD-PV (CDTS) on the KITTI compressed data across multiple code rates, using the PV-RCNN backbone.
+RPKD-PV outperforms SparseKD-PV at all code rates, achieving the highest mAP values.
 Fig.
-5(b) shows the generalization performance of our RPKD framework on the Voxel-RCNN backbone.
-RPKD-Voxel outperforms Voxel-RCNN (baseline) in all code rates, and its mAP value surpasses that of RP-Voxel (T3) and Voxel-RCNN backbone in the high code rate.
-These results vividly illustrate the outstanding robustness and remarkable enhancement capabilities of our method in 3D object detection for compressed point clouds.
-For the generalization ability in WOD, we evaluate the detection performance of our RPKD framework across multiple code rates with the PV-RCNN backbone on the validation split.
-In Table II, RPKD-PV improves mAP/mAPH for PCC-S-C12, PCC-S-C11, and PCC-S-C10 by 0.29/2.1, 0.26/1.77, and 0.86/1.48, respectively.
-For the compressed point clouds of PCC-S-C12, the APH values for vehicle and pedestrian categories exceed those of RP-PV (T2) and PV-RCNN backbone on raw data.
-The experimental results demonstrate that our RPKD framework maintains consistent effectiveness throughout diverse code rates, backbone networks and detection datasets.
+6(b) shows the generalization performance of our RPKD framework on the Voxel-RCNN and SECOND backbones.
+Both RPKD-Voxel (CDTS) and RPKD-SECOND (CDTS) surpass their respective baselines across all code rates, with the two-stage RPKD-Voxel achieving higher mAP than the single-stage RPKD-SECOND.
+These results vividly illustrate the exceptional robustness and significant performance improvements of our method for 3D object detection on compressed point clouds.
+To assess the generalization ability of our RPKD framework, we evaluated its compressed data detection performance with the PV-RCNN backbone on DAIR-V2X-V.
+Table III further shows that RPKD-PV increases mAP by 4.77, 3.34, and 1.28 on the DAIR-V2X-V validation split.
+These experimental results demonstrate that our RPKD framework consistently delivers strong performance across varying code rates, backbone networks, and detection datasets.
 | Options of RKD | Detectors of DKD | mAP | ||||
-| Voxel-wise | Point-wise | PV-RCNN | RP-PV | |||
+| Voxel-wise | Point-wise | w/o RP | w/ RP | |||
 | 71.39 | ||||||
@@
 As shown in Table III, when using the single-layer network with and , the compressed point-cloud detector achieves optimal detection performance.
-| Data Type | Method | Car 3D | Ped. 3D | Cyc. 3D | mAP | ||||||
-| Easy | Mod. | Hard | Easy | Mod. | Hard | Easy | Mod. | Hard | |||
-| Raw Data | PV-RCNN | 91.97 | 84.72 | 82.32 | 68.60 | 60.61 | 55.55 | 92.06 | 74.48 | 69.63 | 75.56 |
-| RP-PV(T2) | 92.10 | 84.36 | 82.29 | 63.62 | 56.35 | 50.88 | 90.30 | 70.75 | 66.96 | 73.07 | |
-| PCC-S-C12 (Bpp=3.864) | PV-RCNN(Baseline) | 91.48 | 82.51 | 80.72 | 56.17 | 49.66 | 45.37 | 87.85 | 68.51 | 64.21 | 69.61 |
-| RPKD-PV(Unified) | 91.46 | 82.92 | 80.01 | 65.55 | 58.14 | 53.16 | 90.54 | 67.30 | 62.98 | 72.45 | |
-| PCC-S-C11 (Bpp=2.146) | PV-RCNN(Baseline) | 90.85 | 82.25 | 80.33 | 57.79 | 50.83 | 46.20 | 88.92 | 65.79 | 61.40 | 69.37 |
-| RPKD-PV(Unified) | 91.20 | 82.37 | 79.29 | 66.71 | 59.32 | 54.18 | 87.66 | 66.86 | 62.26 | 72.21 | |
-| PCC-S-C10 (Bpp=1.047) | PV-RCNN(Baseline) | 90.74 | 81.40 | 79.24 | 54.71 | 47.58 | 44.12 | 81.08 | 61.64 | 57.45 | 66.44 |
-| RPKD-PV(Unified) | 91.37 | 82.37 | 79.48 | 65.38 | 58.05 | 52.91 | 85.69 | 61.66 | 57.44 | 70.48 | |
 IV-C2 Critical Elements of Our RPKD
-As indicated in Table IV, we assessed the influences of critical elements in our RPKD framework on the detection performance for compressed point clouds.
-In the first two rows, we compared the pre-training performance of compressed point-cloud detectors, where refers to the voxel-based averaging operation in the RCM module.
-denotes screening foreground compressed points in calculating reflectance distillation loss.
-In the third row, we used the matched raw point reflectance as the distillation label during point-wise RKD and did not perform the voxel-based .
-In contrast, we utilized the matched reflectance of raw point-cloud voxels as the distillation label during voxel-wise RKD and conducted the voxel-based , as shown in the fourth row.
-In Table IV, PV-RCNN and RP-PV represent raw point-cloud detector types that provided the first-stage proposals for DKD.
-As indicated in the last row, our RPKD framework achieves the best performance with these settings.
+As shown in Table IV, we evaluated the impact of critical elements in our RPKD framework on the compressed data detection performance.
+In the first two rows, we compared the pretraining performance of compressed point-cloud detectors, where refers to the voxel-based averaging operation in the RCM module.
+denotes the process of screening foreground compressed points for RKD.
+In the third row, we used the matched raw point reflectance as the distillation label during point-wise RKD, without performing the voxel-based .
+In contrast, the fourth row employed the matched reflectance of raw point-cloud voxels as the distillation label during voxel-wise RKD, with the voxel-based .
+The sixth and final rows assessed the effect of incorporating the RP mechanism into the teacher model, which provides the first-stage proposals for DKD.
+As indicated in the last row, our RPKD framework achieves the best performance under these settings.
 IV-C3 Designed Modules of Our RPKD
-As shown in Table V, we evaluated the designed module’s effectiveness in 3D object detection for compressed point clouds.
-The first row displays the detection performance of PV-RCNN (baseline) for compressed point clouds of PCC-S-C11.
-We then incrementally introduced the RP module, RKD and DKD into the detector and reported corresponding mAP values.
-All the designed modules improve the performance of PV-RCNN (baseline), with the best mAP value achieved when all modules are utilized together.
-IV-C4 Experiments for a Unified Model
-To develop a unified detector applicable to various code rates, we combined the KITTI training splits of PCC-S-C10, PCC-S-C11, and PCC-S-C12 in a ratio as input for the unified model.
-The trained model was then used to perform inference on each validation split of these compressed point clouds to evaluate detection performance.
-As shown in Table VI, RPKD-PV (unified) increases mAP by 2.84 for both PCC-S-C12 and PCC-S-C11, and by 4.04 for PCC-S-C10.
-These results demonstrate that our RPKD framework under a unified model achieves consistent effectiveness across diverse code rates.
-IV-C5 Visual Comparisons of Detection Results
-As depicted in Fig.
-6, we performed visual comparisons of different data types and detection methods on the KITTI validation split.
-For compressed point-cloud detectors, RPKD-PV achieves fewer false positives and missed detections across multiple code rates, demonstrating the stronger robustness of our RPKD framework in 3D object detection for compressed point clouds.
-However, compared to the raw point-cloud detector based on PV-RCNN, RPKD-PV tends to miss objects with fewer points or at farther distances, which reveals a limitation of our method in detecting such challenging objects.
-IV-C6 Visual Comparisons of Reflectance Prediction
-In each BEV grid, we selected the maximum reflectance value from either raw data or compressed point clouds for coloring the corresponding visual points.
+As shown in Table V, we evaluated the designed module’s effectiveness for 3D object detection on compressed point clouds.
+The first row reports the baseline performance of PV-RCNN (STS-C) on PCC-S-C11.
+We then incrementally integrated the RP module, RKD and DKD into the detector and recorded the corresponding mAP values.
+Each module contributes to performance improvement, with the highest mAP achieved when all modules are combined.
+| Method | Stage | Car 3D | Ped. 3D | Cyc. 3D | Overall |
+| PointPillars (STS-C) [27] | One | 80.45 | 45.22 | 64.76 | 63.48 |
+| CenterPoint (STS-C) [28] | 80.87 | 50.74 | 65.88 | 65.83 | |
+| SECOND (STS-C) [25] | 83.06 | 47.80 | 66.81 | 65.89 | |
+| RPKD-SECOND (CDTS) | 83.72 | 52.89 | 70.90 | 69.17 | |
+| PV-RCNN (STS-C) [23] | Two | 84.48 | 51.61 | 72.03 | 69.37 |
+| PV-RCNN++ (STS-C) [26] | 84.56 | 58.88 | 71.88 | 71.77 | |
+| Voxel-RCNN (STS-C) [24] | 84.97 | 55.19 | 73.41 | 71.19 | |
+| RPKD-PV (CDTS) | 84.76 | 63.38 | 72.66 | 73.60 | |
+| RPKD-Voxel (CDTS) | 84.95 | 62.86 | 74.70 | 74.17 |
+IV-C4 Comparison of Single-Stage and Two-Stage Methods
+In Table VI, we compared our method against advanced single-stage detectors (PointPillars, CenterPoint, and SECOND) and two-stage detectors (PV-RCNN, PV-RCNN++, and Voxel-RCNN) on PCC-S-C11.
+Our RPKD (CDTS) improves overall mAP by 3.28, 4.23, and 2.98 on the SECOND, PV-RCNN, and Voxel-RCNN backbones, respectively.
+Notably, RPKD-Voxel (CDTS) achieves the highest overall mAP across all detection methods.
+These results demonstrate that our method enhances detection performance for both single-stage and two-stage detectors, with two-stage models exhibiting stronger detection robustness.
+IV-C5 Visual Results on DAIR-V2X-V
+Fig.
+7 presents a visual comparison on the DAIR-V2X-V validation split.
+RPKD-PV (CDTS) produces fewer false positives and missed detections across multiple code rates, confirming its robustness.
 As shown in Fig.
-7, the predicted reflectance in compressed point clouds closely matches the actual reflectance in raw data within the major reflective areas of foreground objects.
-These results demonstrate that our RP module accurately predicts the reflectance of foreground objects across various code rates, thereby improving detection robustness for compressed point clouds.
-IV-D Discussion of Limitation and Future Work
-For non-reflectance compressed point clouds, this paper effectively supplements the missing reflectance information through the RP module and RKD, thus enhancing detection robustness.
-However, as shown in Table I and Table II, the APs for pedestrians in WOD and for moderate-level cars and bicycles in KITTI still fall short of the corresponding APs achieved with raw point clouds.
-This may be because RPKD focuses on transferring detection knowledge via DKD without performing explicit geometric enhancement on lossy compressed point clouds.
-In future work, we plan to explore geometric enhancement modules at the receiver to recover missing spatial details, thereby further improving the robustness of 3D object detection against non-reflectance compressed point clouds.
+8, we used the maximum reflectance value from either raw or compressed data to color the corresponding visual points within each BEV grid.
+For PCC-S-C12 and PCC-S-C11, the predicted reflectance from compressed data closely matches the actual reflectance from raw data in the primary reflective regions of foreground objects, highlighting the effectiveness of the RP module.
+However, for PCC-S-C10, the limited number and low density of compressed points lead to a reduction in accuracy, diminishing the reliability of the reflectance prediction.
+IV-D Analysis of Compression Transmission
+In large-scale V2X scenarios, early fusion is constrained by the high bandwidth required to transmit raw LiDAR data.
+To address this, we employed voxel-based lossy geometric compression and omitted reflectance encoding, thereby maximizing bandwidth savings.
+For compression transmission–based cooperative perception, our analysis and evaluation proceed as follows.
+IV-D1 Memory and Bandwidth of Raw LiDAR Data
+Onboard LiDAR sensors capture point cloud data consisting of single-precision floating-point coordinates (X, Y, Z) and 8-bit unsigned integer reflectance values.
+The memory usage of a coordinate and reflectance component is denoted as .
+Accordingly, the memory per raw point is .
+For LiDAR sensors in the KITTI and DAIR-V2X-V datasets, the number of points acquired per second is .
+The corresponding transmission bandwidth is , with the detailed requirements summarized in Table VII.
+| Dataset | Name | Symbol | Value | Unit |
+| Common Parameters | Single Coordinate Memory | 32 | Bit | |
+| Reflectance Memory | 8 | Bit | ||
+| Single Point Memory | 104 | Bit | ||
+| KITTI | Points Per Second | - | ||
+| Transmission Bandwidth | 135.2 | Mbps | ||
+| DAIR-V2X-V | Points Per Second | - | ||
+| Transmission Bandwidth | 74.88 | Mbps |
+In vehicular networking, prior studies [59, 60, 61] report that the channel capacity for connected vehicles typically ranges from 30 Mbps to 1 Gbps under varying traffic conditions.
+Within these limits, transmitting raw point clouds across multiple vehicles overloads both uplink and downlink channels, making real-time data sharing impractical.
+This highlights the necessity of point cloud compression, where data are compressed and encoded before transmission to enable efficient and feasible cooperative perception.
+IV-D2 Bandwidth Calculation for Compressed Point-Cloud Transmission
+Existing LiDAR point cloud compression modes can be classified into three categories based on the degree of information loss: lossless compression, lossy compression, and lossy geometry compression.
+In G-PCC [5], the number of points per second in losslessly compressed point clouds remains identical to that of the raw point clouds, denoted as .
+The corresponding transmission bandwidth is defined as:
+where and denote the code rates for coordinates and reflectance in losslessly compressed point clouds, respectively.
+In G-PCC–based lossy compression, the number of compressed points decreases according to the octree depth.
+Let denote the number of points per second in lossy compressed point clouds, thus the transmission bandwidth is expressed as:
+where and represent the code rates for coordinates and reflectance in lossy compressed point clouds, respectively.
+Moreover, lossy geometry compression further reduces transmission bandwidth by discarding reflectance encoding.
+Assuming the number of points per second remains , the transmission bandwidth for lossy geometry compressed point clouds is obtained by the following formula:
+IV-D3 Evaluation for Multi-Vehicle Communication
+Based on G-PCC, Table VIII reports the multi-vehicle communication performance of different compression modes on the KITTI dataset.
+The key metrics include coordinate Bpp, reflectance Bpp, transmission bandwidth, and channel margin.
+Since coordinate and reflectance code rates lack a fixed proportional relationship, we selected representative values from prior compression studies [9, 7] under medium compression ratio settings, corresponding to a point retention rate of 0.85.
+The channel capacity for connected vehicles was set to 200 Mbps, a level achievable in most traffic scenarios.
+The channel margin for a given vehicle is defined as:
+where and denote the bandwidth required for receiving and sending point clouds with a single neighboring vehicle, and is the number of connected vehicles, all adopting the same compression setting.
+| Compression Mode | Point Retention Rate | Points Per Second | Coordinate Bpp | Reflectance Bpp | Bandwidth (Mbps) | Channel Capacity (Mbps) | Channel Margin (Mbps) | |
+| Uncompressed | 1.0 | - | - | 135.2 | 200 | -70.4 | -2504.0 | |
+| Lossless | 1.0 | 20.22 | 4.76 | 32.47 | 200 | 135.06 | -449.4 | |
+| Lossy | 0.85 | 3.81 | 1.68 | 6.04 | 200 | 187.92 | 79.2 | |
+| Lossy Geometry | 0.85 | 3.81 | - | 4.19 | 200 | 191.62 | 116.2 | |
+As shown in Table VIII, lossless, lossy, and lossy geometry compression reduce transmission bandwidth by factors of 3, 21, and 31, respectively.
+For two-vehicle communication (), the corresponding channel margin ratios are 67.5%, 94.0%, and 95.8%.
+These results confirm that point cloud compression effectively alleviates bandwidth demands, with lossy geometry compression achieving the lowest bandwidth and highest margin.
+When scaled to larger communication scenarios (), the channel margin ratios for lossy and lossy geometry compression remain at 39.6% and 58.1%, respectively, underscoring the advantage of discarding reflectance encoding in large-scale cooperative perception.
+IV-D4 Evaluation for Compression Transmission on DAIR-V2X-V
+In real V2X scenarios, we assessed the effectiveness of lossy geometry compression and transmission on the DAIR-V2X-V dataset, using metrics such as compression ratio, coordinate Bpp, and transmission bandwidth.
+The LiDAR sensor in this dataset operates with 40 beams at a 10 Hz capture frequency, producing approximately points per second.
+As reported in Table IX, PCC-S-C12, PCC-S-C11, and PCC-S-C10 achieve compression ratios of 24.55, 45.39, and 98.74, respectively, reducing transmission bandwidth by factors of 34, 94, and 373.
+These findings demonstrate that the PCC-S lossy geometry compression substantially reduces transmission costs in practical V2X environments, enabling real-time sharing of LiDAR data.
+| Data Type | Points Per Second | Compression Ratio | Coordinate Bpp | Bandwidth (Mbps) |
+| Raw Data | - | - | 74.88 | |
+| PCC-S-C12 | 24.55 | 4.01 | 2.17 | |
+| PCC-S-C11 | 45.39 | 2.13 | 0.79 | |
+| PCC-S-C10 | 98.74 | 0.96 | 0.2 |
+IV-E Discussion of Limitation and Future Work
+For non-reflectance compressed data, we supplemented the missing reflectance information using the RP module and RKD, thereby enhancing detection robustness.
+Nevertheless, as shown in Table I, the APs for moderate-level cars and bicycles in KITTI still lag behind those achieved on raw data.
+A likely reason is that our RPKD framework emphasizes transferring detection knowledge via DKD without explicitly enhancing the geometry of lossy compressed point clouds.
+Moreover, this study represents an initial step toward compression transmission-based collaborative perception, with the benchmark task limited to single-view 3D object detection.
+In real-world collaborative scenarios, connected vehicles will need to integrate locally collected raw data with decoded compressed data to expand perception range and improve detection accuracy.
+Therefore, effectively leveraging heterogeneous data of varying quality for registration, fusion, and 3D object detection remains a critical open challenge.
+In future work, we plan to develop geometric enhancement modules at the receiver to recover missing spatial details, further strengthening the robustness of 3D object detection on non-reflectance compressed data.
+Beyond this, we aim to advance collaborative 3D object detection by jointly exploiting raw and enhanced compressed data, thereby pushing forward the development of compression transmission-based collaborative perception.
 V Conclusion
 We have presented a 3D object detection framework with reflectance prediction-based knowledge distillation (RPKD) to improve detection accuracy for non-reflectance compressed point clouds.
-During detection training, the designed RCM module assigns the average reflectance of raw point-cloud voxels to corresponding compressed points based on their spatial relationship, providing essential prior labels for reflectance reconstruction.
-Treating non-reflectance compressed point clouds as weak students, the geometry-based RP module plays a critical role in precise reflectance prediction and object detection within compressed point-cloud detectors.
-The raw point-cloud detector, with the same structure, transfers the accurate reflectance prediction and object detection knowledge through RKD and DKD, significantly enhancing the robustness of compressed point-cloud detectors.
-Experimental results on the KITTI dataset and WOD demonstrate that our RPKD framework consistently improves detection performance for compressed point clouds across various code rates and backbone networks.
+During training, the RCM module assigns the average reflectance of raw point-cloud voxels to corresponding compressed points based on their spatial relationships, providing essential prior labels for reflectance reconstruction.
+By treating non-reflectance compressed point clouds as weak data, the geometry-based RP module plays a critical role in accurate reflectance prediction and object detection within compressed point-cloud detectors.
+Building on CDTS, the raw point-cloud detector with the same structure transfers distillation knowledge through RKD and DKD, significantly boosting the robustness of compressed point-cloud detectors.
+Experimental results on the KITTI and DAIR-V2X-V datasets demonstrate that our RPKD framework consistently improves detection performance across various code rates and backbone networks.